### v1

# Accelerating Natural Gradient Descent for PINNs with Randomized Numerical Linear Algebra

Abstract
Natural Gradient Descent (NGD) has emerged as a promising optimization algorithm for training neural network-based solvers for partial differential equations (PDEs), such as Physics-Informed Neural Networks (PINNs). However, its practical use is often limited by the high computational cost of solving linear systems involving the Gramian matrix. While matrix-free NGD methods based on the conjugate gradient (CG) method avoid explicit matrix inversion, the ill-conditioning of the Gramian significantly slows the convergence of CG. In this work, we extend matrix-free NGD to broader classes of problems than previously considered and propose the use of Randomized Nyström preconditioning to accelerate convergence of the inner CG solver. The resulting algorithm demonstrates substantial performance improvements over existing NGD-based methods on a range of PDE problems discretized using neural networks.
1 Introduction
Neural network-based approaches for solving partial differential equations (PDEs) have gained significant attention as promising alternatives to traditional numerical methods for both forward and inverse problems. A common strategy is to embed the governing PDE directly into the loss function, as in Physics-Informed Neural Networks (PINNs) [24, 59, 63], the Deep Ritz method [25], and related variants [41, 7, 32, 8, 22]. These methods hold the promise of providing mesh-free approximations of PDE solutions and of exploiting the dimension-independent approximation power of neural networks [16]. However, their practical adoption is often hindered by optimization difficulties. First-order methods like Adam [42] typically perform poorly, plateauing at low accuracy [46, 67, 60, 68]. This limitation is generally attributed not to the expressiveness of neural networks, but to the ill-conditioning of the loss [60, 21, 46]. Consequently, specialized optimization strategies are increasingly being advocated [19, 55, 62].
More principled second-order optimization algorithms have only recently begun to emerge [54, 55, 21, 73, 20, 60, 33, 38]. Among them, (Energy) Natural Gradient Descent (NGD) [54] has shown remarkable performance, achieving high accuracy in few iterations by using a search direction of the form $\mathbf{d}=-(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})^{-1}\nabla L(\boldsymbol{\theta})$, where $\boldsymbol{\theta}\in\mathbb{R}^{p}$ denotes the neural network parameters, $L$ is the loss function, $\mathbf{G}(\boldsymbol{\theta})$ the positive semidefinite Gramian matrix (see Section 2.2), and $\mu>0$ a regularization parameter. However, solving linear systems with $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$ using direct methods has a cubic cost $\mathcal{O}(p^{3})$, making it impractical for large-scale networks. To mitigate this, matrix-free strategies have been proposed [38, 73], where the system is solved approximately using the conjugate gradient (CG) method using only matrix-vector products (matvecs) with $\mathbf{G}(\boldsymbol{\theta})$, without assembling the matrix. In practice, however, $\mathbf{G}(\boldsymbol{\theta})$ is often ill-conditioned, as shown in Figure 1, leading to slow CG convergence and motivating the need for preconditioning. In this work, we address this issue using tools from Randomized Numerical Linear Algebra (RandNLA) [51, 53]. Our main contributions can be summarized as follows:
  • We demonstrate how to efficiently compute matrix-vector products with the Gramian in a more general setting than those considered in [55, 38, 73].
  • We propose the use of randomized Nyström preconditioning [26] to construct a preconditioner for $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$ via a low-rank approximation of the Gramian. This turns the strong spectral decay of the Gramian from a challenge into an asset for our approach.
  • We introduce a novel algorithm, NyströmNGD, that combines matrix-free NGD with Randomized Nyström preconditioning.
  • We present results showing significant improvements over prior NGD-based methods.
Related works
Various strategies have been proposed to address training pathologies in PINNs and other neural network-based PDE solvers. These include adaptive quadrature point selection and loss term weighting [67, 13, 66, 69, 71, 68]. More recently, second-order optimizers derived from preconditioning the underlying infinite-dimensional PDE operator have shown promise [21, 55, 54, 33]. Natural Gradient Descent (NGD) has emerged as a unifying framework for these methods, as each can be seen as NGD under a different choice of the optimization metric [56, 55]. Matrix-free implementations have been explored for specific metrics [38, 73], but none of these works addresses the ill-conditioning of the Gramian, which hampers CG convergence in practice. To address this bottleneck, our work bridges NGD with Randomized Numerical Linear Algebra (RandNLA). The use of RandNLA in Machine Learning is well-established [27, 23, 65]. Randomized Nyström preconditioning [26] has been applied to PINNs in [60], but only as a final stage after Adam and L-BFGS, and applying the Nyström approximation, designed for symmetric positive semidefinite matrices, to the Hessian, which may be indefinite. Alternative notions of preconditioning, where the preconditioner is embedded into the loss function, are explored in [47, 7]. K-FAC has also been used to approximate the Gramian [20], but not as a preconditioner, and it requires problem-specific adaptations.
Figure 1: Training of a PINN for the 3D Poisson problem with exact solution $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$, using Energy Natural Gradient Descent [54]. Left: spectral decay of the Gramian $\mathbf{G}(\boldsymbol{\theta})$ at different iterations. Right: $\mathrm{H}^{1}(\Omega)$ relative error during training.
2 Preliminaries
2.1 Notation
Let $\mathrm{V}$ be a real Hilbert space. Its inner product and norm are denoted by $\left({\cdot},\,{\cdot}\right)_{\mathrm{V}}$ and $\left\|{\cdot}\right\|_{\mathrm{V}}$. Its (topological) dual is indicated $\mathrm{V}^{\prime}$, and the duality pairing by $\left\langle{\cdot},\,{\cdot}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}$. The space of bounded linear operators from $\mathrm{V}$ to another Hilbert space $\mathrm{W}$ is denoted by $\mathcal{L}(\mathrm{V},\mathrm{W})$. We say that $\mathcal{A}\in\mathcal{L}(\mathrm{V},\mathrm{V}^{\prime})$ is symmetric positive semidefinite (SPSD) if the associated bilinear form $a(u,v)=\left\langle{\mathcal{A}u},\,{v}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}$ is symmetric and satisfies $a(u,u)\geq 0$ for all $u\in\mathrm{V}$. It is symmetric positive definite (SPD) if $a$ is also coercive, i.e., if $\exists\,\alpha>0$ such that $a(u,u)\geq\alpha\left\|{u}\right\|_{\mathrm{V}}^{2}$ for all $u\in\mathrm{V}$.
Let $\left(\mathcal{M},g\right)$ be a smooth Riemannian manifold. The tangent space at $u\in\mathcal{M}$ is denoted by $\mathrm{T}_{u}\mathcal{M}$, with inner product and norm given by ${g_{u}}\left({\cdot},\,{\cdot}\right)$ and $\left\|{\cdot}\right\|_{g_{u}}$, respectively. For a smooth cost function $E:\mathcal{M}\to\mathbb{R}$, the Riemannian gradient and Hessian are denoted by $\mathrm{grad}^{g}E$ and $\mathrm{Hess}^{g}E$, respectively. The Euclidean gradient of a function $E:\mathbb{R}^{n}\to\mathbb{R}$ is denoted by $\nabla E$.
2.2 Natural Gradient Descent
Consider the minimization problem on a Riemannian manifold $(\mathcal{M},g)$
$\inf_{u\in\mathcal{M}}E(u).$ (1)
To approximate its solution, we introduce a parametrized ansatz $\mathcal{N}=\left\{u_{\boldsymbol{\theta}}\coloneqq\rho(\boldsymbol{\theta})\,:\,\boldsymbol{\theta}\in\Theta\right\}\subseteq\mathcal{M}$, where $\rho:\Theta\to\mathcal{M}$ is a smooth parametrization map. We then consider the optimization problem
$\inf_{\boldsymbol{\theta}\in\Theta}L(\boldsymbol{\theta})\coloneqq\inf_{\boldsymbol{\theta}\in\Theta}E(u_{\boldsymbol{\theta}})=\inf_{\boldsymbol{\theta}\in\Theta}(E\circ\rho)(\boldsymbol{\theta}).$
For instance, $u_{\boldsymbol{\theta}}$ could be a neural network with parameters $\boldsymbol{\theta}$, though the framework applies more generally. We will refer to $\Theta\subseteq\mathbb{R}^{p}$ as the parameter space, and to $\mathcal{M}$ as the function space.
Natural Gradient Descent (NGD) [3, 50, 56, 55] follows the steepest descent direction according to the metric on $\mathcal{M}$, the “natural” one for the function $u_{\boldsymbol{\theta}}$, and pulls it back to the parameter space. Let $\boldsymbol{\theta}$ be the current iterate. The steepest descent direction in function space would be $-\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}})$. In NGD, the corresponding the parameter update direction is $-\mathbf{d}$, where the natural gradient $\mathbf{d}\in\mathbb{R}^{p}$ is such that the first-order expansion $u_{\boldsymbol{\theta}-\alpha\mathbf{d}}\approx u_{\boldsymbol{\theta}}-\alpha\mathrm{D}\rho(\boldsymbol{\theta})[\mathbf{d}]$ best approximates (in the metric $g$) the update in function space. This corresponds to $\mathrm{D}\rho(\boldsymbol{\theta})[\mathbf{d}]=\sum_{j=1}^{p}d_{j}\cdot\partial_{\theta_{j}}u_{\boldsymbol{\theta}}$ being the $g_{u_{\boldsymbol{\theta}}}$-orthogonal projection of $\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}})$ onto the model’s generalized tangent space $\mathrm{T}_{\boldsymbol{\theta}}\mathcal{N}=\mathrm{Span}\left(\left\{\partial_{\theta_{1}}u_{\boldsymbol{\theta}},\dots,\partial_{\theta_{p}}u_{\boldsymbol{\theta}}\right\}\right)\subset\mathrm{T}_{u_{\boldsymbol{\theta}}}\mathcal{M}.$ As shown, e.g, in [3, 56], the natural gradient is given by $\mathbf{d}=\mathbf{G}(\boldsymbol{\theta})^{\dagger}\nabla_{\boldsymbol{\theta}}L(\boldsymbol{\theta})$, where $\mathbf{G}(\boldsymbol{\theta})^{\dagger}$ is the Moore–Penrose pseudoinverse of the matrix
$\left[\mathbf{G}(\boldsymbol{\theta})\right]_{ij}={g_{u_{\boldsymbol{\theta}}}}\left({\partial_{\theta_{i}}u_{\boldsymbol{\theta}}},\,{\partial_{\theta_{j}}u_{\boldsymbol{\theta}}}\right),$ (2)
often referred to as the Gramian [72, 55] or Fisher information matrix [3, 50]. As observed in the literature [72, 12], it is often beneficial to add a damping term to improve numerical stability and convergence, yielding the regularized direction
$\mathbf{d}=\left(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I}\right)^{-1}\nabla_{\boldsymbol{\theta}}L(\boldsymbol{\theta}),$ (3)
where $\mu>0$ is a regularization parameter, typically adapted during the optimization process.
2.3 Randomized Nyström preconditioning
The Gramian matrix $\mathbf{G}=\mathbf{G}(\boldsymbol{\theta})$ (2) is positive semidefinite and often numerically rank-deficient, as shown in Figure 1. In this setting, Randomized Nyström preconditioning [26] enables efficient construction of a preconditioner for $(\mathbf{G}+\mu\mathbf{I})$ using only matrix-vector products with $\mathbf{G}$. This preconditioner will be used with conjugate gradient to compute the regularized natural gradient (3).
Randomized Nyström approximation
Let $\mathbf{G}\in\mathbb{R}^{p\times p}$ be symmetric positive semidefinite (PSD). The randomized Nyström approximation [51, Section 14] with a random test matrix $\boldsymbol{\Omega}\in\mathbb{R}^{p\times\ell}$ is
$\hat{\mathbf{G}}_{\mathrm{nys}}\coloneqq(\mathbf{G}\boldsymbol{\Omega})\left(\boldsymbol{\Omega}^{\top}\mathbf{G}\boldsymbol{\Omega}\right)^{\dagger}(\mathbf{G}\boldsymbol{\Omega})^{\top}.$ (4)
The randomness in $\boldsymbol{\Omega}$ ensures that $\hat{\mathbf{G}}_{\mathrm{nys}}$ approximates $\mathbf{G}$ well with high probability [51, Section 14.4]. In this work, we take $\boldsymbol{\Omega}$ to be standard normal. This choice, over sampling-based methods [2, 5] and other alternatives [4, 51, 53, 31], is motivated by stronger theoretical guarantees [26, 51] and by the fact that in our setting $\mathbf{G}=\mathbf{G}(\boldsymbol{\theta})$ in (11) is accessible only via matrix-vector products, and therefore structured test matrices requiring entrywise access are not applicable or less advantageous.
We emphasize that formula (4) is numerically unstable and should not be used directly. A more stable and efficient implementation, which also returns the eigendecomposition of $\hat{\mathbf{G}}_{\mathrm{nys}}$, is described in [51, Algorithm 16] and included in Appendix A.
Randomized Nyström preconditioner
Given an eigendecomposition $\hat{\mathbf{G}}_{\mathrm{nys}}=\mathbf{U}\hat{\boldsymbol{\Lambda}}\mathbf{U}^{\top}$ with $\hat{\boldsymbol{\Lambda}}=\mathrm{diag}(\hat{\lambda}_{1},\dots,\hat{\lambda}_{\ell})$ and $\hat{\lambda}_{1}\geq\dots\geq\hat{\lambda}_{\ell}$, the randomized Nyström preconditioner [26] for $(\mathbf{G}+\mu\mathbf{I})$ with $\mu\geq 0$ is
$\mathbf{P}^{-1}=(\hat{\lambda}_{\ell}+\mu)\mathbf{U}(\hat{\boldsymbol{\Lambda}}+\mu\mathbf{I})^{-1}\mathbf{U}^{\top}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\top}).$ (5)
To ensure good performance in expectation, the sampling parameter $\ell$ should be proportional to the effective dimension $d_{\mathrm{eff}}(\mu)=\sum_{i=1}^{p}\lambda_{i}(\mathbf{G})/(\lambda_{i}(\mathbf{G})+\mu)$, which acts as a smoothed count of eigenvalues of $\mathbf{G}$ larger than $\mu$ [26, Theorem 1.1]. When $\mathbf{G}$ exhibits strong spectral decay, $d_{\mathrm{eff}}(\mu)\ll p$, making the randomized Nyström preconditioner particularly effective.
Computational complexity
Each application of the randomized Nyström preconditioner costs $\mathcal{O}(p\ell)$ flops. Beforehand, one must compute the Nyström approximation. Denoting by $\mathrm{C}_{\mathrm{mv}}$ the cost of a matrix-vector product with $\mathbf{G}$, the total cost is $\mathcal{O}(\mathrm{C}_{\mathrm{mv}}\ell+p\ell^{2})$ flops, with memory requirement $\mathcal{O}(p\ell)$. The term $\mathrm{C}_{\mathrm{mv}}\ell$ comes from computing $\mathbf{G}\boldsymbol{\Omega}$, i.e., $\ell$ independent matvecs with $\mathbf{G}$, and is thus embarrassingly parallel. This allows for significant acceleration on multi-core CPUs or GPUs.
3 Matrix-free Natural Gradient Descent
Prior work has shown that matrix-vector products with the Gramian $\mathbf{G}(\boldsymbol{\theta})$ (2) can be computed via automatic differentiation for specific metrics [38, 55, 73]. We present a unified framework for metric selection and show that autodiff-based matvecs extend to broader metric classes.
3.1 Choice of the Metric
In some cases, the problem’s structure naturally suggests a metric, as in Variational MonteCarlo [58, 36, 55] or Sobolev gradient flows [35, 17]. In general, the link between NGD and Riemannian gradient descent [54] suggests selecting the metric $g$ to reduce the condition number of the Riemannian Hessian [14]. Further details are deferred to Appendix B.
Linear problems
Consider a linear problem $\mathcal{A}u=f\in\mathrm{W}^{\prime}$, where $\mathrm{V},\mathrm{W}$ are Hilbert spaces, and $\mathcal{A}:\mathrm{V}\to\mathrm{W}^{\prime}$ is a linear isomorphism. The associated least-squares problem and natural metric are
$\inf_{u\in\mathrm{V}}E(u)\coloneqq\frac{1}{2}\left\|{\mathcal{A}u-f}\right\|_{\mathrm{W}^{\prime}}^{2},\qquad{g}\left({v},\,{w}\right)\coloneqq\left({\mathcal{A}v},\,{\mathcal{A}w}\right)_{\mathrm{W}^{\prime}}.$ (6)
If $\mathrm{V}=\mathrm{W}$ and $\mathcal{A}$ is symmetric positive definite, one can endow $\mathrm{V}$ with the $\mathcal{A}$-induced inner product $\left({v},\,{w}\right)_{\mathcal{A}}=\left\langle{\mathcal{A}v},\,{w}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}$, and use it as the NGD metric. Then, up to an additive constant irrelevant for optimization, the loss (6) coincides with the variational form used in the Deep Ritz method [25]
$\inf_{u\in\mathrm{V}}E(u)\coloneqq\frac{1}{2}\left\langle{\mathcal{A}u},\,{u}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}-\left\langle{f},\,{u}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}.$ (7)
A link with operator preconditioning
NGD can be viewed as a preconditioned gradient descent, where the preconditioner is the Gramian [21]. For a linear parametrized model $u_{\boldsymbol{\theta}}=\sum_{i=1}^{p}\theta_{i}\,\varphi_{i}$, the Gramian becomes $\left[\mathbf{G}(\boldsymbol{\theta})\right]_{ij}={g_{u_{\boldsymbol{\theta}}}}\left({\varphi_{i}},\,{\varphi_{j}}\right)$. If $(\mathcal{M},g)=(\mathrm{V},\left({\cdot},\,{\cdot}\right)_{\mathrm{V}})$ and $g$ is induced by a positive definite linear operator $\mathcal{G}:\mathrm{V}\to\mathrm{V}^{\prime}$, then $\mathbf{G}(\boldsymbol{\theta})=\mathbf{G}$ where $[\mathbf{G}]_{ij}=\left\langle{\mathcal{G}\varphi_{i}},\,{\varphi_{j}}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}$ does not depend on $\boldsymbol{\theta}$. This coincides with preconditioners arising from an operator preconditioning perspective, in the context of Galerkin discretizations of PDEs [37, 48, 43, 30].
Nonlinear problems
If $\mathcal{M}$ is equipped with a metric $a$ and $\mathrm{Hess}^{a}E(u)$ is positive definite for all $u\in\mathcal{M}$, Newton’s method [14] is induced by choosing the metric
${g_{u}}\left({v},\,{w}\right)=\mathrm{Hess}^{a}E(u)[v,w].$ (8)
In the case $(\mathcal{M},a)=(\mathrm{V},\left({\cdot},\,{\cdot}\right)_{\mathrm{V}})$ for a Hilbert space $\mathrm{V}$, this leads to the Energy Natural Gradient Descent (ENGD) method [54, 55]. Note that to interpret ENGD as a Natural Gradient Descent, it is necessary that $\mathrm{Hess}^{a}E(u)$ is SPD for all $u\in\mathcal{M}$, as otherwise (8) does not define a valid metric.
Consider now the nonlinear problem $\mathcal{R}(u)=0$, where $\mathcal{R}:\mathrm{V}\to\mathrm{W}$ is a (possibly nonlinear) residual between Hilbert spaces $\mathrm{V},\mathrm{W}$. For the least-squares functional $E(u)=\frac{1}{2}\left\|{\mathcal{R}(u)}\right\|_{\mathrm{W}}^{2}$, a symmetric positive semidefinite approximation of the Hessian is the Gauss-Newton one [1, 70]:
${g_{u}}\left({v},\,{w}\right)=\left({\mathrm{D}\mathcal{R}(u)[v]},\,{\mathrm{D}\mathcal{R}(u)[w]}\right)_{\mathrm{W}}.$ (9)
This choice of the metric leads the Gauss-Newton NGD method [38, 33]. Note that $g_{u}$ is, in general, only semidefinite. Positive definitness of $g_{u}$ is necessary to interpret it as a Riemannian metric, and whether this condition holds depends on the structure of the underlying nonlinear PDE.
3.2 Matrix-free formulation
We show that matrix-vector products with the Gramian (2) can be computed via automatic differentiation, provided the metric satisfies the following assumption.
Assumption 3.1.
Let $\Omega\subset\mathbb{R}^{d}$ be a domain, and let $\mathcal{M}\subseteq\mathrm{V}$, where $\mathrm{V}$ is a Hilbert space of functions on $\Omega$. Assume that the metric $g_{u}$ takes the form
${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\mathcal{F}_{u}v(\mathbf{x})\cdot\mathcal{F}_{u}w(\mathbf{x})\,\varphi_{u}(\mathbf{x})\,\mathrm{d}\mathbf{x},$ (10)
for all $v,w\in\mathrm{T}_{u}\mathcal{M}\subseteq\mathrm{V}$, where $\mathcal{F}_{u}:\mathrm{V}\to\mathrm{L}^{2}(\Omega;\varphi)^{k}$ is linear and continuous.
Theorem 3.2.
Let 3.1 hold, and let $\mathtt{sg}:\mathbb{R}^{p}\to\mathbb{R}^{p}$ denote the stop-gradient operator, formally defined by $\mathtt{sg}(\boldsymbol{\theta})=\boldsymbol{\theta}$ and $\nabla\mathtt{sg}(\boldsymbol{\theta})=\mathbf{0}$. If the integral (10) is approximated using a quadrature rule with points $\{\mathbf{x}_{r}\}_{r=1}^{q}$ and weights $\{w_{r}\}_{r=1}^{q}$, then
$\begin{split}\mathbf{G}(\boldsymbol{\theta})&=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}(\boldsymbol{\theta})^{\top}\begin{bmatrix}w_{1}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{1})\mathbf{I}_{k}&&\\ &\ddots&\\ &&w_{q}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{q})\mathbf{I}_{k}\end{bmatrix}\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}(\boldsymbol{\theta}),\quad\,\,\,\mathbf{F}(\boldsymbol{\theta})=\begin{bmatrix}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{1})\\ \vdots\\ \mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{q})\end{bmatrix},\end{split}$ (11)
where $\mathbf{J}_{\boldsymbol{\theta}}$ denotes the Jacobian with respect to $\boldsymbol{\theta}$.
The proof of 3.2 is deferred to Appendix C. As a consequence of 3.2, matrix-vector products with $\mathbf{G}(\boldsymbol{\theta})$ can be computed using only products with the Jacobian and the transpose Jacobian of $\mathbf{F}$. These operations can be efficiently implemented via automatic differentiation, using forward-mode autodiff for Jacobian-vector products (JVPs) and backward-mode autodiff for vector-Jacobian products (VJPs). Note that most autodiff frameworks include a stop-gradient function.
Remark 3.3.
3.1 encompasses standard metrics used in PDE-related function spaces. For instance, the $\mathrm{L}^{2}(\Omega)$ and $\mathrm{H}^{1}(\Omega)$ metrics correspond to the choices $\mathcal{F}_{u}v=v$ and $\mathcal{F}_{u}v=(\nabla v,v)$, respectively. It also includes the Gauss-Newton metric (9) on $\mathrm{L}^{2}(\Omega)$, obtained by taking $\mathcal{F}_{u}=\mathrm{D}\mathcal{R}(u)$, for which matrix-free NGD implementations have been proposed in [38, 55] for specific problems. However, 3.1 is significantly more general and covers a much broader class of metrics. In practice, it includes many of the metrics relevant in the context of PDEs, since these typically need to be expressible as integrals to be computationally tractable. For example, it includes the Newton metric ${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\nabla v\cdot\nabla w+3u^{2}vw$ in [55, Section 4.2], and the energy inner product for the Gross–Pitaevskii equation as considered in [35], ${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\nabla v\cdot\nabla w+Vvw+\beta\lvert u\rvert^{2}vw$.
Computational Cost
With automatic differentiation, the cost of each JVP and VJP is a small constant factor (typically between $2$ and $4$) times the cost of evaluating $\mathbf{F}_{i}(\boldsymbol{\theta})$ for $i=1,2$ [29, 49]. For a multilayer perceptron, the evaluation cost scales as $\mathcal{O}(pq)$, where $p$ is the number of trainable parameters and $q$ is the number of quadrature points. Therefore, the cost of each matrix-vector product with $\mathbf{G}(\boldsymbol{\theta})$ is $\mathrm{C}_{\mathrm{mv}}=\mathcal{O}(pq)$. Note however that when executed on modern GPUs, batch evaluation across $q$ points benefits from parallelization.
4 NyströmNGD: Matrix-Free Natural Gradient Descent with Randomized Nyström Preconditioning
To compute the regularized natural gradient, we approximately solve $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})\mathbf{d}=\nabla L(\boldsymbol{\theta})$ using the preconditioned conjugate gradient (PCG) method in a matrix-free fashion, with the preconditioner constructed via randomized Nyström. This approach leverages the structure of the problem: matrix-vector products with the Gramian $\mathbf{G}(\boldsymbol{\theta})$ can be computed efficiently via automatic differentiation, while the strong spectral decay commonly observed in practice (see Figure 1) permits an accurate low-rank approximation, yielding a preconditioner that is both efficient to apply and effective in accelerating convergence. This section presents the resulting algorithm in full detail.
Choice of the regularization parameter $\mu$
The regularization parameter $\mu>0$ is crucial for numerical stability and for mitigating the slow initial convergence typical of quasi-second-order methods like NGD. Prior works proposed heuristics that adapt $\mu$ dynamically [12, 28] or decrease it with the loss [72], so that early iterations mimic gradient descent and gradually shift to NGD near a local minimum. We instead propose a heuristic based on the condition number of $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$.
Letting $\lambda_{1}\geq\dots\geq\lambda_{p}$ denote the eigenvalues of the Gramian $\mathbf{G}(\boldsymbol{\theta})$, the spectral condition number of $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$ is $k=(\lambda_{1}+\mu)/(\lambda_{p}+\mu)\leq 1+\lambda_{1}/\mu$. Hence, choosing $\mu=c\cdot\lambda_{1}$ guarantees $k\leq 1+c^{-1}$. In practice, good results are obtained by setting $c$ proportional to machine epsilon, i.e., $\mu=\gamma\cdot\epsilon_{\mathrm{mach}}\cdot\lambda_{1}$, with $\gamma=p$ as a reasonable default. Observe that although $\lambda_{1}$ is not directly available, it can be estimated by $\hat{\lambda}_{1}$ from randomized Nyström. It can be shown that the expected relative error can be small even for modest values of $\ell$, when $\mathbf{G}(\boldsymbol{\theta})$ exhibits strong spectral decay [51, Theorem 14.1].
Choice of rank parameter $\ell$
The rank parameter $\ell$ controls the computational cost of the preconditioner and should ideally scale with the effective dimension $d_{\mathrm{eff}}(\mu)$ [26, Theorem 1.1]. As $d_{\mathrm{eff}}(\mu)$ is not directly accessible, we estimate $\ell$ adaptively. Following [26, Section 5.4.2], we do it by monitoring the ratio $\hat{\lambda}_{\ell}/\mu$: if it drops below $10$, we double $\ell$ for the next iteration, up to a prescribed maximum $\ell_{\max}$. Otherwise, we reduce $\ell$ to the smallest index satisfying $\hat{\lambda}_{\ell}/\mu<10$, adding a small offset (e.g., $1$ or $2$) to avoid oscillations.
Other algorithmic details and pseudocode
Algorithm 1 summarizes the proposed algorithm. At each iteration, we construct the randomized Nyström preconditioner (lines 5 to 7) using the matrix-free Gramian. We then apply preconditioned conjugate gradient (PCG) to solve the linear system (3) (line 9). The PCG stopping criterion combines a maximum number of iterations maxit and a relative tolerance $\mathtt{rel\_tol}=\min(\kappa,\left\|{\nabla L(\boldsymbol{\theta}_{k})}\right\|_{2})$, inspired by the Steihaug–Toint truncated CG method [18]. This design avoids overspending computation far from critical points, where NGD steps may underperform compared to GD, and ensures accuracy near local minima. Once the search direction $\mathbf{d}_{k}$ is computed, a line search is performed to update the iterate (lines 10 and 11). In practice, we found that a standard backtracking line-search [70] suffices, without requiring logarithmic grid searches as in [54, 38]. Finally, the rank parameter $\ell$ is updated adaptively (lines 12 to 16).
1:Initial parameter $\boldsymbol{\theta}_{0}\in\mathbb{R}^{p}$, initial rank $\ell_{0}$, maximum rank $\ell_{\max}$, $\gamma\in\mathbb{R}_{+}$ (e.g. $\gamma=p$), CG parameters $\mathtt{maxit},\kappa$ (e.g. $\mathtt{maxit}=20,\kappa=0.1$)
2:Initialize $\ell=\ell_{0}$
3:for $k=0,1,2,\dots$ do
4:     Define $\mathbf{G}_{\mathrm{fun}}=\mathbf{v}\mapsto\mathbf{G}(\boldsymbol{\theta})\mathbf{v}$ via (11) $\triangleright$ Matrix-free Gramian
5:     $[\mathbf{U},\hat{\boldsymbol{\Lambda}}]=\mathtt{RandomizedNystrom}(\mathbf{G}_{\mathrm{fun}},\ell)$triangleright Randomized Nyström approximation
6:     $\mu=\gamma\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}}$triangleright Regularization parameter $\mu$
7:     $\mathbf{P}_{\mathrm{fun}}=\mathbf{v}\mapsto\frac{1}{\hat{\lambda}_{\ell}+\mu}\mathbf{U}(\hat{\boldsymbol{\Lambda}}+\mu\mathbf{I})\mathbf{U}^{\top}\mathbf{v}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\top})\mathbf{v}$triangleright Matrix-free preconditioner
8:     $\mathtt{rel\_tol}=\min(\kappa,\left\|{\nabla L(\boldsymbol{\theta}_{k})}\right\|_{2})$triangleright Relative tolerance for PCG
9:     $\mathbf{d}_{k}=\texttt{PCG}(\mathbf{G}_{\mathrm{fun}}+\mu\cdot\mathrm{Id},\,\nabla L(\boldsymbol{\theta}_{k}),\,\mathtt{rel\_tol},\,\mathtt{maxit},\,\mathbf{P}_{\mathrm{fun}})$triangleright Run PCG
10:     $\alpha_{k}=\texttt{Linesearch}(\boldsymbol{\theta}_{k},-\mathbf{d}_{k})$triangleright Linesearch along $\alpha\mapsto\boldsymbol{\theta}_{k}-\alpha\mathbf{d}_{k}$
11:     $\boldsymbol{\theta}_{k+1}=\boldsymbol{\theta}_{k}-\alpha_{k}\mathbf{d}_{k}$
12:     if $\hat{\lambda}_{\ell}<10\mu$ then
13:         $\ell=\min(2\ell,\ell_{\max})$triangleright Possibly increase rank parameter $\ell$
14:     else
15:         $\ell=\min\left(\min\{i:\hat{\lambda}_{i}>10\mu\}+1,\,\ell_{\max}\right)$triangleright Possibly decrease rank parameter $\ell$
16:     end if
17:end for
Algorithm 1 NyströmNGD
5 Numerical Experiments
We test NyströmNGD on a range of forward and inverse PDE problems with known analytical solutions. Additional resources and implementation details are provided in Appendix E. Additional numerical experiments involving variational formulations of PDEs are presented in Appendix F.
Comparison of the methods
We compare NyströmNGD against Natural Gradient Descent with SVD-based pseudoinverse (legend: NGD), matrix-free NGD solved via conjugate gradient without preconditioning (legend: NGD with CG), and the BFGS algorithm [70]. We run NGD-based methods for $500$ or $1000$ iterations, depending on the computational cost, and BFGS for $6\times 10^{4}$ or $1.2\times 10^{5}$ iterations. For a fair comparison with NyströmNGD, in NGD with CG we use the same relative tolerance as in Algorithm 1 and set the maximum number of iterations to $\texttt{maxit}+\ell_{\max}$. Given the different per-iteration cost of the methods, we compare them in both iteration count and wall-clock time. We do not include first-order optimizers (e.g., gradient descent, Adam) or L-BFGS, as they have been shown to be non-competitive [38, 55]. In all experiments, we employ fully connected neural networks with $\tanh$ activation and two hidden layers of width $32$. Integrals in both the loss and the metric are discretized via Monte Carlo quadrature. All experiments are averaged over 10 runs. In the plots, the solid line denotes the median, and the shaded region spans the first and third quantiles.
Computational details
For the PINN formulations, we use JAX [15] and JAXopt [11]. The JAXopt BFGS implementation was optimized by us for improved efficiency, as detailed in Appendix D, yielding, toghether with JIT compilation, a 1000×1000$\times$1000 × speedup over the timings reported in [54]. The NGD implementation is taken from the public repository of [72]. All experiments were conducted in double precision on a single NVIDIA RTX 4090 GPU with 24GB of memory.
5.1 Poisson problem in 3D
We consider the Poisson problem on the 3D unit cube
$\begin{cases}-\Delta u=f&\Omega=(0,1)^{3},\\ u=g_{D}&\partial\Omega,\end{cases}$ (12)
with exact solution $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$ as in [72]. The PINNs least-squares loss and the corresponding metric are
$E(u)=\frac{1}{2}\left\|{\Delta u+f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(\partial\Omega)}^{2},$
${g}\left({u},\,{v}\right)=\left({\Delta u},\,{\Delta v}\right)_{\mathrm{L}^{2}(\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\partial\Omega)}=\frac{1}{2}\int_{\Omega}\Delta u\,\Delta v\,\mathrm{d}\mathbf{x}+\frac{1}{2}\int_{\partial\Omega}u\,v\,\mathrm{d}\mathbf{s}.$
Further details are provided in Section E.1.
As shown in Figure 2, NyströmNGD yields the best results in both final accuracy and computational efficiency, converging in around $100$ iterations. The Nyström rank parameter $\ell$ increases rapidly during early iterations, consistent with Figure 1, and then stabilizes well below the maximum allowed rank $\ell_{\max}=500$. This confirms the effectiveness of the adaptive rank selection. NGD initially converges faster in iterations but plateaus early, while NGD with CG is slower due to the many CG steps needed to approximate the Gramian inverse. BFGS is not competitive in iteration count, though its lower per-iteration cost allows for faster initial progress in wall-clock time. However, it ultimately stagnates at a higher error.
Figure 2: Poisson equation in 3D discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $500$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
5.2 Heat equation in 3+1D
We consider the heat equation
$\begin{cases}\partial_{t}u-\Delta u=f&I\times\Omega=(0,1)\times(0,1)^{3},\\ u=g_{D}&I\times\partial\Omega,\\ u(0,\cdot)=u_{0}&\Omega,\end{cases}$ (13)
with exact solution $u(t,x,y,z)=\left(\cos(\pi x)+\cos(\pi y)+\cos(\pi z)\right)e^{-\pi^{2}t/4}$ as in [72]. The corresponding least squares loss and metrics are
$E(u)=\frac{1}{2}\left\|{\partial_{t}u-\Delta u-f}\right\|_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\partial\Omega))}^{2}+\frac{1}{2}\left\|{u(0,\cdot)-u_{0}}\right\|_{\mathrm{L}^{2}(\Omega)}^{2},$
${g}\left({u},\,{v}\right)=\left({\partial_{t}u-\Delta u},\,{\partial_{t}v-\Delta v}\right)_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}+\left({u(0,\cdot)},\,{v(0,\cdot)}\right)_{\mathrm{L}^{2}(\Omega)}.$
Further details are provided in Section E.2.
Also in this example, NyströmNGD provides the best accuracy at a reduced computational cost, as shown in Figure 3. All other methods saturate at a lower accuracy, despite being given a higher computational budget. The Nyström rank parameter $\ell$ rapidly increases and then stabilizes at a value lower than $\ell_{\max}=500$, although at a higher value compared to the Poisson example.
Figure 3: Heat equation in 3+1D discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
5.2.1 Kovasznay Flow (2D steady state Navier–Stokes flow)
We consider the two-dimensional steady Navier–Stokes flow as originally described in [45], with Reynolds number $\mathrm{Re}=40$, following the setup in [38, Section 4.1]. The system of PDEs reads
$\begin{cases}-\nu\Delta u+(u\cdot\nabla)u+\nabla p=f&\Omega=(-\frac{1}{2},1)\times(-\frac{1}{2},\frac{3}{2}),\\ \nabla\cdot u=0&\Omega,\\ u=g_{D}&\partial\Omega.\end{cases}$ (14)
The nonlinear least squares loss is
$E(u,p)=\frac{1}{2}\left\|{-\nu\Delta u+(u\cdot\nabla)u+\nabla p-f}\right\|^{2}_{\mathrm{L}^{2}(\Omega)^{2}}+\frac{1}{2}\left\|{\nabla\cdot u}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{1}{2}\left\|{u-g_{D}}\right\|^{2}_{\mathrm{L}^{2}(\partial\Omega)},$
and as a metric we consider the Gauss–Newton metric (9). We employ two separate neural networks for $u$ and $p$. Further details are provided in Section E.3.
NyströmNGD still provides the best accuracy in fewer than $1000$ iterations, as shown in Figure 4. However, compared to previous examples, the rank $\ell$ grows to the maximum rank $\ell_{\max}=500$ before stabilizing around $400$, indicating that the Gramian is less suited to low-rank approximation than in previous examples. This explains the smaller difference, both in final accuracy and computation time, between NyströmNGD and BFGS. Nevertheless, the gap in iterations remains substantial.
Figure 4: Kovasnay Flow discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
5.3 Source recovery problem for the Poisson equation in 3D
We consider the inverse problem of recovering both the solution $u$ and the source term $f$ in the 3D Poisson equation (12) from $N_{\mathrm{data}}=40$ noiseless observations $\{u^{\star}_{i}\}_{i=1}^{N_{\mathrm{data}}}$ of the true solution at random points $\{\mathbf{x}_{i}\}_{i=1}^{N_{\mathrm{data}}}$ in $\Omega$. The corresponding least squares loss and metrics are
$E(u,f)=\frac{1}{2}\left\|{\Delta u+f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(\partial\Omega)}^{2}+\frac{1}{2N_{\mathrm{data}}}\sum_{i=1}^{N_{\mathrm{data}}}\lvert u(\mathbf{x}_{i})-u^{\star}_{i}\rvert^{2}$
${g}\left({(u,f)},\,{(v,h)}\right)=\left({\Delta u+f},\,{\Delta v+h}\right)_{\mathrm{L}^{2}(\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\partial\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\Omega)}.$
We employ two separate neural networks for $u$ and $f$. Further details are provided in Section E.4.
As shown in Figure 5, NyströmNGD achieves higher accuracy than NGD and NGD with CG, and converges in fewer iterations than BFGS. However, BFGS can perform more iterations within the same time budget and eventually surpasses NyströmNGD in accuracy. The reduced performance of NyströmNGD can be explained by the poor low-rank approximability of the Gramian, as indicated by the rank parameter $\ell$ reaching the threshold $\ell_{\max}=500$. This suggests that $\ell_{\max}$ matrix-vector products are insufficient for accurately approximating the Gramian via randomized Nyström, leading to a less effective preconditioner.
Figure 5: Source recovery problem for the Poisson equation in 3D discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations (first $500$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
6 Discussion and Conclusion
We introduced NyströmNGD, a novel optimization algorithm that uses randomized Nyström preconditioning to accelerate the inner conjugate gradient (CG) solver in matrix-free natural gradient descent (NGD). This approach significantly reduces the cost of the inner solver, resulting in faster convergence in both iterations and wall-clock time. Moreover, our work extends matrix-free NGD to a broader class of metrics relevant to PDE-related problems. We validated our method on different PDE problems discretized using PINNs, demonstrating that NyströmNGD consistently outperforms prior NGD methods by achieving higher accuracy at a lower computational cost.
Limitations and future work
The use of randomized Nyström for preconditioning builds on the empirically observed low-rank structure of the Gramian. Our results confirm that in many cases the Gramian admits strong low-rank compression, though this is not always guaranteed and warrants further analysis. Nevertheless, the core idea of preconditioning the inner CG solver remains applicable, and we plan to explore alternative Gramian approximations for use as preconditioners. Notably, NyströmNGD not only reduces the computational cost but also improves final accuracy over standard NGD. This may be due to improved numerical stability or the stochastic nature of the Nyström approximation, which could help escape poor local minima, similar to how SGD can outperform full-batch GD in traditional machine learning [44, 39, 34]. This aspect merits further investigation. Finally, NyströmNGD is especially promising in the context of neural network-based solvers for variational formulations of PDEs, where the choice of metric can be guided by the underlying variational structure. Our preliminary results (see Appendix F) support the potential of the proposed approach and motivate future research in this direction.
Acknowledgments and Disclosure of Funding
This work has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie Action MSCA-DN-101119556 (IN-DEEP). The authors are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). C. Marcati and G. Sangalli acknowledge support from the Italian MUR through the PRIN 2022 PNRR project NOTES (No. P2022NC97R). G. Sangalli further acknowledges support from the Italian MUR through PNRR-M4C2-I1.4-NC-HPC-Spoke6. C. Marcati further acknowledges the PRIN 2022 project ASTICE (No. 202292JW3F).
Appendix A Pseudocode of Randomized Nyström Approximation
1:PSD matrix $\mathbf{G}\in\mathbb{R}^{p\times p}$ (as a function $\mathbf{v}\mapsto\mathbf{G}\mathbf{v}$), rank $\ell$
2:Nyström approximation $\hat{\mathbf{G}}_{\mathrm{nys}}=\mathbf{U}\hat{\boldsymbol{\Lambda}}\mathbf{U}^{\top}$
3:$\boldsymbol{\Omega}=\textrm{randn}(p,\ell)$triangleright Gaussian test matrix
[…]
5:$\mathbf{Y}=\mathbf{G}\boldsymbol{\Omega}$triangleright $\ell$ matvecs with $\mathbf{G}$
[…]
7:$\mathbf{Y}_{\nu}=\mathbf{Y}+\nu\boldsymbol{\Omega}$triangleright Shift for stability
8:$\mathbf{C}=\textrm{chol}(\boldsymbol{\Omega}^{\top}\mathbf{Y}_{\nu})$
9:$\mathbf{B}=Y_{\nu}/\mathbf{C}$triangleright Triangular solve
10:$[\mathbf{U},\boldsymbol{\Sigma},\sim]=\textrm{svd}(\mathbf{B},0)$triangleright Thin SVD
11:$\hat{\boldsymbol{\Lambda}}=\max\{0,\boldsymbol{\Sigma}^{2}-\nu\mathbf{I}\}$triangleright Remove shift
Algorithm 2 Randomized Nyström Approximation [26, 51]
Appendix B An extended discussion on the choice of the metric
In this Appendix section, we further elaborate on the selection of the metric discussed in Section 3.1, providing additional details and extending the discussion.
It can be shown that NGD corresponds to Riemannian gradient descent [14] on the manifold $(\mathcal{M},g)$, up to the $g$-orthogonal projection of the Riemannian gradient onto the model’s generalized tangent space $\mathrm{T}_{\boldsymbol{\theta}}\mathcal{N}$, and a second-order correction; see [55, Theorem 1]. More precisely, the NGD update $\boldsymbol{\theta}_{k+1}=\boldsymbol{\theta}_{k}-\alpha_{k}\mathbf{G}(\boldsymbol{\theta}_{k})^{\dagger}\nabla L(\boldsymbol{\theta}_{k})$ satisfies
$u_{\boldsymbol{\theta}_{k+1}}=u_{\boldsymbol{\theta}_{k}}-\alpha_{k}\Pi_{\boldsymbol{\theta}_{k}}^{g}\left(\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}_{k}})\right)+\varepsilon_{k},$
where $\Pi_{\boldsymbol{\theta}_{k}}^{g}$ denotes the $g$-orthogonal projection onto the generalized tangent space $\mathrm{T}_{\boldsymbol{\theta}_{k}}\mathcal{N}$, and $\varepsilon_{k}=\mathcal{O}\big{(}\alpha_{k}^{2}\left\|{\mathbf{G}(\boldsymbol{\theta}_{k})^{\dagger}\nabla L(\boldsymbol{\theta}_{k})}\right\|^{2}\big{)}$. A principled strategy for selecting the metric $g$ is thus to adopt criteria from Riemannian optimization, choosing $g$ to reduce the condition number of the Riemannian Hessian [14, 1]. Although convergence may differ from standard Riemannian gradient descent due to the projection onto $\mathrm{T}_{\boldsymbol{\theta}}\mathcal{N}$, whose effect depends on the geometry of $\mathcal{N}$ and the parametrization, this strategy can be expected to remain effective if the ansatz $\mathcal{N}$ is sufficiently expressive.
B.1 Linear problems
Setting B.1.
Let $\mathrm{V},\mathrm{W}$ be Hilbert spaces, and let $\mathcal{A}:\mathrm{V}\to\mathrm{W}^{\prime}$ be a linear isomorphism. Given $f\in\mathrm{W}^{\prime}$ we aim at solving
$\mathcal{A}u=f\in\mathrm{W}^{\prime}.$ (15)
A natural approach is to consider the least-squares problem
$\inf_{u\in\mathrm{V}}E(u)\coloneqq\frac{1}{2}\left\|{\mathcal{A}u-f}\right\|_{\mathrm{W}^{\prime}}^{2}.$ (16)
Choosing on $\mathcal{M}=\mathrm{V}$ the (point-independent) metric
${g}\left({v},\,{w}\right)\coloneqq\left({\mathcal{A}v},\,{\mathcal{A}w}\right)_{\mathrm{W}^{\prime}},$ (17)
yields $\mathrm{Hess}^{g}E=\mathrm{Id}$. Riemannian gradient descent on $(\mathrm{V},g)$ thus converges in a single iteration, as the Riemannian gradient is given by $\mathrm{grad}^{g}E(u)=u-\mathcal{A}^{-1}f$, i.e., computing it is equivalent to solving (15). NGD does not require applying $\mathcal{A}^{-1}$: computing the Gramian $\left[\mathbf{G}(\boldsymbol{\theta})\right]_{ij}=\left({\mathcal{A}\partial_{\theta_{i}}u_{\boldsymbol{\theta}}},\,{\mathcal{A}\partial_{\theta_{j}}u_{\boldsymbol{\theta}}}\right)_{\mathrm{W}^{\prime}}$ involves only applications of $\mathcal{A}$.
Remark B.2.
We formulate the linear problem from $\mathrm{V}$ to the dual of $\mathrm{W}$ because one of our target applications involves variational formulations of PDEs, which naturally take the form (15). For PINNs, which are based on strong formulations, $\mathrm{W}$ is typically a (product of) $\mathrm{L}^{2}$ spaces, so that $\mathrm{W}^{\prime}\cong\mathrm{W}$ and the formulation recovers that of [72], which does not employ the dual.
Remark B.3.
In Setting B.1, it is typically straightforward to approximate numerically the inner product $\left({\cdot},\,{\cdot}\right)_{\mathrm{W}}$ and the duality pairing $\left\langle{\cdot},\,{\cdot}\right\rangle_{\mathrm{W}^{\prime}\times\mathrm{W}}$, but not the dual inner product $\left({\cdot},\,{\cdot}\right)_{\mathrm{W}^{\prime}}$. Let $\mathcal{R}_{\mathrm{W}}:\mathrm{W}\to\mathrm{W}^{\prime}$ denote the Riesz map, defined by
$\left\langle{\mathcal{R}_{\mathrm{W}}v},\,{w}\right\rangle_{\mathrm{W}^{\prime}\times\mathrm{W}}=\left({v},\,{w}\right)_{\mathrm{W}}\qquad\forall v,w\in\mathrm{W}.$
Then, for $\ell,h\in\mathrm{W}^{\prime}$, we have
$\left({\ell},\,{h}\right)_{\mathrm{W}^{\prime}}=\left\langle{\ell},\,{\mathcal{R}_{\mathrm{W}}^{-1}h}\right\rangle_{\mathrm{W}^{\prime}\times\mathrm{W}}.$
For example, when $\mathrm{W}=\mathrm{H}^{1}_{0}(\Omega)$, the dual space is $\mathrm{W}^{\prime}=\mathrm{H}^{-1}(\Omega)$, and $\mathcal{R}_{\mathrm{W}}=\Delta$, the Laplace operator. Computing the dual inner product thus requires applying $(-\Delta^{-1})$, i.e., solving a (discretized) Laplace problem. In contrast, for $\mathrm{W}=\mathrm{L}^{2}(\Omega)$, one has $\mathrm{W}^{\prime}\cong\mathrm{W}$ and $\mathcal{R}_{\mathrm{W}}=\mathrm{Id}$.
Symmetric positive definite linear operators
If $\mathrm{V}=\mathrm{W}$ and $\mathcal{A}$ is symmetric positive definite, one can endow $\mathrm{V}$ with the $\mathcal{A}$-induced inner product
$\left({v},\,{w}\right)_{\mathcal{A}}=\left\langle{\mathcal{A}v},\,{w}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}\qquad\forall v,w\in\mathrm{V}.$
This corresponds to choosing $\mathcal{R}_{\mathrm{V}}=\mathcal{A}$, and the loss (16) becomes
$E(u)=\frac{1}{2}\left\langle{\mathcal{A}u},\,{u}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}-\left\langle{f},\,{u}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}+\mathrm{const},$ (18)
where the constant is irrelevant for gradient-based optimization. This is precisely the variational form used in the Deep Ritz method [25]. The corresponding natural metric is $g=\left({\cdot},\,{\cdot}\right)_{\mathcal{A}}$.
B.2 Nonlinear problems
For general nonlinear problems, there is no canonical choice of metric as in the linear case. In some instances, the structure of the problem suggests multiple metrics; the choice of the metric impacts the convergence and the computational cost of the optimization schemes. Examples include Variational Monte Carlo (VMC) [58, 36, 55] or Sobolev gradient flows [35, 17]. Below we outline some general strategies that can be used in nonlinear problems and are employed in our numerical experiments.
Newton’s method
Setting B.4.
Let $(\mathcal{M},a)$ be a Riemannian manifold and $f:\mathcal{M}\to\mathbb{R}$ a smooth function. We aim to solve $\inf_{u\in\mathcal{M}}E(u)$.
If $\mathrm{Hess}^{a}E(u)$ is symmetric positive definite for all $u\in\mathcal{M}$, a natural metric leading to Riemannian Newton’s method [14] is
${g_{u}}\left({v},\,{w}\right)=\mathrm{Hess}^{a}E(u)[v,w].$ (19)
In the case $(\mathcal{M},a)=(\mathrm{V},\left({\cdot},\,{\cdot}\right)_{\mathrm{V}})$ for a Hilbert space $\mathrm{V}$, this yields the Energy Natural Gradient Descent (ENGD) method [54, 55]. Note that to interpret ENGD as a Natural Gradient Descent, it is necessary that $\mathrm{Hess}^{a}E(u)$ is SPD for all $u\in\mathcal{M}$, as otherwise (19) does not define a valid metric.
Nonlinear least-squares problem: Gauss-Newton
Setting B.5.
Let $\mathrm{V},\mathrm{W}$ be Hilbert spaces, and let $\mathcal{R}:\mathrm{V}\to\mathrm{W}$ be a (possibly nonlinear) residual. We aim at finding $u\in\mathrm{V}$ such that $\mathcal{R}(u)=0$.
A natural approach is to minimize the nonlinear least-squares functional
$\inf_{u\in\mathrm{V}}E(u)\coloneqq\inf_{u\in\mathrm{V}}\frac{1}{2}\left\|{\mathcal{R}(u)}\right\|_{\mathrm{W}}^{2}.$ (20)
The Hessian of $f$ reads
$\mathrm{D}^{2}E(u)[v,w]=\left({\mathrm{D}\mathcal{R}(u)[v]},\,{\mathrm{D}\mathcal{R}(u)[w]}\right)_{\mathrm{W}}+\left({\mathcal{R}(u)},\,{\mathrm{D}^{2}\mathcal{R}(u)[v,w]}\right)_{\mathrm{W}},$
and a symmetric positive semidefinite approximation is given by the Gauss-Newton method [70]
${g_{u}}\left({v},\,{w}\right)=\left({\mathrm{D}\mathcal{R}(u)[v]},\,{\mathrm{D}\mathcal{R}(u)[w]}\right)_{\mathrm{W}}.$ (21)
This choice of the metric leads the Gauss-Newton NGD method [38, 33]. Note that $g_{u}$ is, in general, only semidefinite. While this does not pose numerical issues when using the pseudoinverse or damping, the positive definitness of $g_{u}$ is necessary to interpret it as a Riemannian metric. Whether this condition holds depends on the structure of the underlying nonlinear PDE.
Appendix C Proof of 3.2
See 3.1
See 3.2
Proof.
Let $\mathbf{e}_{h}\in\mathbb{R}^{p}$ denote the $h$-th vector of the canonical basis, i.e., the vector that has a $1$ in the $h$-th entry and 00 in all other entries. Using the linearity and continuity of $\mathcal{F}_{u_{\boldsymbol{\theta}}}$,
$\begin{split}\mathcal{F}_{u_{\boldsymbol{\theta}}}\partial_{\theta_{h}}u_{\boldsymbol{\theta}}&=\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}\partial_{\theta_{h}}u_{\boldsymbol{\theta}}=\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}^{k}\left(\lim_{t\to 0}\frac{u_{\boldsymbol{\theta}+t\mathbf{e}_{h}}-u_{\boldsymbol{\theta}}}{t}\right)\\ &=\lim_{t\to 0}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}\left(\frac{u_{\boldsymbol{\theta}+t\mathbf{e}_{h}}-u_{\boldsymbol{\theta}}}{t}\right)=\lim_{t\to 0}\frac{\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}+t\mathbf{e}_{h}}-\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}}{t}\\ &=\partial_{\theta_{h}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}.\end{split}$
Note that the use of the stop-gradient operator is crucial to ensure that we do not differentiate $\mathcal{F}_{u_{\boldsymbol{\theta}}}$ with respect to $\boldsymbol{\theta}$. Hence,
$\begin{split}[\mathbf{G}(\boldsymbol{\theta})]_{ij}&={g_{u_{\boldsymbol{\theta}}}}\left({\partial_{\theta_{i}}u_{\boldsymbol{\theta}}},\,{\partial_{\theta_{j}}u_{\boldsymbol{\theta}}}\right)=\int_{\Omega}\mathcal{F}_{u_{\boldsymbol{\theta}}}\partial_{\theta_{i}}u_{\boldsymbol{\theta}}\cdot\mathcal{F}_{u_{\boldsymbol{\theta}}}\partial_{\theta_{j}}u_{\boldsymbol{\theta}}(\mathbf{x})\,\varphi_{u_{\boldsymbol{\theta}}}\,\mathrm{d}\mathbf{x}\\ &=\int_{\Omega}\partial_{\theta_{i}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}\cdot\partial_{\theta_{j}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}\,\varphi_{u_{\boldsymbol{\theta}}}\,\mathrm{d}\mathbf{x},\end{split}$
or, in matrix form,
$\mathbf{G}(\boldsymbol{\theta})=\left(\mathrm{D}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}\right)^{*}\,\mathrm{D}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}},$
where the Hilbert adjoint $\left(\mathrm{D}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}\right)^{*}$ is taken with respect to the inner product on $\mathrm{L}^{2}(\Omega,\varphi)^{k}$, and we identify a matrix with the induced linear map.
Applying quadrature with nodes $\{\mathbf{x}_{r}\}_{r=1}^{q}$ and weights $\{w_{r}\}_{r=1}^{q}$, we get
$\begin{split}[\mathbf{G}(\boldsymbol{\theta})]_{ij}&=\sum_{r=1}^{q}w_{r}\,\partial_{\theta_{i}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{r})\cdot\partial_{\theta_{j}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{r})\,\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{r})\end{split}$
and thus, in matrix form:
$\begin{split}\mathbf{G}(\boldsymbol{\theta})&=\sum_{r=1}^{q}w_{r}\,\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{r})\,\left(\mathbf{J}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{r})\right)^{\top}\mathbf{J}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{r})\\ &=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}(\boldsymbol{\theta})^{\top}\begin{bmatrix}w_{1}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{1})\mathbf{I}_{k}&&\\ &\ddots&\\ &&w_{q}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{q})\mathbf{I}_{k}\end{bmatrix}\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}(\boldsymbol{\theta}).\end{split}$
Remark C.1.
More in general, it can be shown that if the metric $g_{u}$ takes the form
${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\mathcal{F}_{u}^{1}v(\mathbf{x})\cdot\mathcal{F}_{u}^{2}w(\mathbf{x})\,\varphi_{u}(\mathbf{x})\,\mathrm{d}\mathbf{x},$ (22)
where $\mathcal{F}_{u}^{i}:\mathrm{V}\to\mathrm{L}^{2}(\Omega;\varphi)^{k}$ are linear and continuous for $i=1,2$, and the integral is approximated using a quadrature rule with points $\{\mathbf{x}_{r}\}_{r=1}^{q}$ and weights $\{w_{r}\}_{r=1}^{q}$, then
$\begin{split}\mathbf{G}(\boldsymbol{\theta})&=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}_{1}(\boldsymbol{\theta})^{\top}\begin{bmatrix}w_{1}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{1})\mathbf{I}_{k}&&\\ &\ddots&\\ &&w_{q}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{q})\mathbf{I}_{k}\end{bmatrix}\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}_{2}(\boldsymbol{\theta}),\quad\,\,\,\mathbf{F}_{i}(\boldsymbol{\theta})=\begin{bmatrix}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}^{i}u_{\boldsymbol{\theta}}(\mathbf{x}_{1})\\ \vdots\\ \mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}^{i}u_{\boldsymbol{\theta}}(\mathbf{x}_{q})\end{bmatrix}.\end{split}$ (23)
Appendix D Modification of jaxopt.BFGS
We refer to the JAXopt implementation of the BFGS algorithm in https://github.com/google/jaxopt/blob/main/jaxopt/_src/bfgs.py (commit c818497). Using the notation from the implementation, lines 240-250 compute the BFGS inverse Hessian update [70] as:
$\mathbf{H}_{k+1}=(\mathbf{I}-\rho\mathbf{s}\mathbf{y}^{\top})\mathbf{H}_{k}(\mathbf{I}-\rho\mathbf{y}\mathbf{s}^{\top})+\rho\mathbf{s}\mathbf{s}^{\top},\quad\rho=\frac{1}{\mathbf{s}^{\top}\mathbf{y}}.$
This involves outer products $\mathbf{s}\mathbf{y}^{\top}$ and $\mathbf{s}\mathbf{s}^{\top}$ and two matrix-matrix multiplications. We replace it with a more efficient formulation that avoids full matrix-matrix products, using vector-matrix operations and rank-one updates:
$\mathbf{H}_{k+1}=\mathbf{H}_{k}+\left[\rho+\rho^{2}\mathbf{y}^{\top}(\mathbf{H}_{k}\mathbf{y})\right]\mathbf{s}\mathbf{s}^{\top}-\rho\left[(\mathbf{H}_{k}\mathbf{y})\mathbf{s}^{\top}+\mathbf{s}(\mathbf{H}_{k}\mathbf{y})^{\top}\right].$
This reformulation eliminates matrix-matrix multiplications, resulting in significant speedups while maintaining numerical stability and equivalence to the original update.
Appendix E Additional resources for the numerical experiments in Section 5
We begin with PDEs in strong form discretized via Physics-Informed Neural Networks (PINNs) [59, 19]. In all PINNs experiments, we set in Algorithm 1 the parameters $\ell_{0}=10$, $\ell_{\max}=500$, $\gamma=p$, $\mathtt{maxit}=20$, and $\kappa=0.1$. For NGD and NGD with CG, we use $\mu=\min(10^{-5},L(\boldsymbol{\theta}))$ as suggested in [72], from which most experiments are adapted. To enforce stronger regularization in early iterations, in NyströmNGD we take the regularization parameter as $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-4}\cdot L(\boldsymbol{\theta})^{2})$. Errors are always computed using an order of magnitude more quadrature points than in training, sampled independently. Random integration points sampled once and fixed throughout training.
E.1 Poisson problem in 3D
We consider the Poisson problem on the 3D unit cube
$\begin{cases}-\Delta u=f&\Omega=(0,1)^{3},\\ u=g_{D}&\partial\Omega,\end{cases}$ (24)
with exact solution $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$ as in [72]. Problem (24) is linear and we consider it in strong form. Using the notation of Section B.1, the function spaces are therefore $\mathrm{V}=\mathrm{H}^{2}(\Omega)$ and $\mathrm{W}=\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\partial\Omega)\cong\mathrm{W}^{\prime}$, with operator
$\mathcal{A}:\mathrm{H}^{2}(\Omega)\to\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\partial\Omega),\qquad\mathcal{A}u=(-\Delta u,\left.{u}\right|_{\partial\Omega}).$
The corresponding least-squares loss is
$E(u)=\frac{1}{2}\left\|{\Delta u+f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(\partial\Omega)}^{2},$ (25)
and the induced metric reads
${g}\left({u},\,{v}\right)=\left({\Delta u},\,{\Delta v}\right)_{\mathrm{L}^{2}(\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\partial\Omega)}=\frac{1}{2}\int_{\Omega}\Delta u\,\Delta v\,\mathrm{d}\mathbf{x}+\frac{1}{2}\int_{\partial\Omega}u\,v\,\mathrm{d}\mathbf{s}.$ (26)
To discretize problem (24) using a PINN formulation, we take $u=u_{\boldsymbol{\theta}}$ a neural network and the loss function $L(\boldsymbol{\theta})=E(u_{\boldsymbol{\theta}})$ is approximated using MonteCarlo quadrature.
Remark E.1.
The metric (26) is coercive only in the $\mathrm{H}^{1/2}(\Omega)$ norm, but not in $\mathrm{H}^{2}(\Omega)$. This is due due to the use of the $\mathrm{L}^{2}(\partial\Omega)$ norm for the Dirichlet boundary term. Coercivity in $\mathrm{H}^{2}(\Omega)$ can be restored by enforcing the boundary conditions strongly [72, Remark 6], or by penalizing in the $\mathrm{H}^{1/2}(\partial\Omega)$ norm [13]. However, in line with standard PINNs practice, we retain the $\mathrm{L}^{2}(\partial\Omega)$ penalty for simplicity.
Computational details
We use a fully connected neural network $u_{\boldsymbol{\theta}}$ with $\tanh$ activation and two hidden layers of width $32$. Integrals in the loss and metric are discretized via Monte Carlo quadrature with $N_{\Omega}=3000$ points uniformly sampled in the interior $\Omega$ and $N_{\partial\Omega}=800$ on the boundary $\partial\Omega$. We run NGD-based methods for $500$ iterations and BFGS for $6\times 10^{4}$ iterations.
Figure 6: Poisson equation in 3D discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $500$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 7: Poisson equation in 3D discretized via PINNs and trained with NGD using SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
E.2 Heat equation in 3+1D
We consider the heat equation
$\begin{cases}\partial_{t}u-\Delta u=f&I\times\Omega=(0,1)\times(0,1)^{3},\\ u=g_{D}&I\times\partial\Omega,\\ u(0,\cdot)=u_{0}&\Omega,\end{cases}$ (27)
with exact solution $u(t,x,y,z)=\left(\cos(\pi x)+\cos(\pi y)+\cos(\pi z)\right)e^{-\pi^{2}t/4}$ as in [72]. We consider Equation 27 in strong form, with function spaces $\mathrm{V}=\mathrm{L}^{2}(I,\mathrm{H}^{2}(\Omega))\cap\mathrm{H}^{1}(I,\mathrm{L}^{2}(\Omega))$, $\mathrm{W}=\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))\times\mathrm{L}^{2}(I,\mathrm{L}^{2}(\partial\Omega))\times\mathrm{L}^{2}(\Omega)$, and linear operator
$\mathcal{A}u=\big{(}\partial_{t}u-\Delta u,\;\left.{u}\right|_{I\times\partial\Omega},\;u(0,\cdot)\big{)}.$
The corresponding least squares loss is
$E(u)=\frac{1}{2}\left\|{\partial_{t}u-\Delta u-f}\right\|_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\partial\Omega))}^{2}+\frac{1}{2}\left\|{u(0,\cdot)-u_{0}}\right\|_{\mathrm{L}^{2}(\Omega)}^{2},$ (28)
with induced metric
${g}\left({u},\,{v}\right)=\left({\partial_{t}u-\Delta u},\,{\partial_{t}v-\Delta v}\right)_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}+\left({u(0,\cdot)},\,{v(0,\cdot)}\right)_{\mathrm{L}^{2}(\Omega)}.$ (29)
Considerations on the coercivity of the metric similar to Remark E.1 apply here. For details, we refer to [72, Section 4.5]. To discretize problem (27) we take $u=u_{\boldsymbol{\theta}}$ a neural network and approximate $L(\boldsymbol{\theta})=E(u_{\boldsymbol{\theta}})$ using MonteCarlo quadrature.
Computational details
We use a fully connected neural network $u_{\boldsymbol{\theta}}$ with $\tanh$ activation and two hidden layers of width $32$. Integrals in the loss and metric are discretized via Monte Carlo quadrature with $N_{I\times\Omega}=3000$ points uniformly sampled in the interior $I\times\Omega$ and $N_{\Gamma}=800$ on the spatio-temporal boundary $\Gamma=(I\times\partial\Omega)\cup(\{0\}\times\Omega)$. We run NGD-based methods for $1000$ iterations and BFGS for $6\times 10^{4}$ iterations.
Figure 8: Heat equation in 3+1D discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 9: Heat equation in 3+1D discretized via PINNs and trained with NGD using SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
E.3 Kovasznay Flow (2D steady state Navier–Stokes flow)
We consider the two-dimensional steady Navier–Stokes flow as originally described in [45], with Reynolds number $\mathrm{Re}=40$, following the setup in [38, Section 4.1]. The system of PDEs reads
$\begin{cases}-\nu\Delta u+(u\cdot\nabla)u+\nabla p=f&\Omega=(-\frac{1}{2},1)\times(-\frac{1}{2},\frac{3}{2}),\\ \nabla\cdot u=0&\Omega,\\ u=g_{D}&\partial\Omega,\end{cases}$ (30)
with exact solution
$\begin{split}u(x,y)&=\begin{bmatrix}1-e^{\lambda x}\cos(2\pi y)\\ \frac{\lambda}{2\pi}e^{\lambda x}\sin(2\pi y)\end{bmatrix},\\ p(x,y)&=\frac{1}{2}(1-e^{2\lambda x})-c,\end{split}$
where $\lambda=\frac{1}{2\nu}-\sqrt{\frac{1}{4\nu^{2}}+4\pi^{2}}$, $\nu=\frac{1}{\mathrm{Re}}=\frac{1}{40}$, and $c$ is chosen such that $\int_{\Omega}p=0$. Following [38], we consider the function spaces $\mathrm{V}=\mathrm{H}^{2}(\Omega)\times\{p\in\mathrm{H}^{1}(\Omega):\int_{\Omega}p=0\}$ and $\mathrm{W}=\mathrm{L}^{2}(\Omega)^{2}\times\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\partial\Omega)\cong\mathrm{W}^{\prime}$. We write problem (30) in strong form as a nonlinear least squares problem with residual
$\mathcal{R}(u,p)=\begin{bmatrix}-\nu\Delta u+(u\cdot\nabla)u+\nabla p-f\\ \nabla\cdot u\\ u-g_{D}\end{bmatrix}$ (31)
and corresponding energy functional
$E(u,p)=\frac{1}{2}\left\|{-\nu\Delta u+(u\cdot\nabla)u+\nabla p-f}\right\|^{2}_{\mathrm{L}^{2}(\Omega)^{2}}+\frac{1}{2}\left\|{\nabla\cdot u}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{1}{2}\left\|{u-g_{D}}\right\|^{2}_{\mathrm{L}^{2}(\partial\Omega)}.$ (32)
As a metric we consider the Gauss–Newton metric (9). We refer to [38] for a discussion on coercivity. To discretize (30), we take $(u,p)=(u_{\boldsymbol{\theta}_{1}},p_{\boldsymbol{\theta}_{2}})$ two neural networks with weights combined into a single vector $\boldsymbol{\theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}]$, and approximate $L(\boldsymbol{\theta})=E(u_{{\boldsymbol{\theta}}_{1}},p_{{\boldsymbol{\theta}}_{2}})$ using Monte Carlo quadrature.
Computational details
We employ fully connected neural networks $u_{\boldsymbol{\theta}_{1}}$ and $p_{\boldsymbol{\theta}_{2}}$ with $\tanh$ activation and two hidden layers of width $32$. Integrals in the loss and metric are discretized via Monte Carlo quadrature with $N_{\Omega}=3000$ points uniformly sampled in the interior $\Omega$ and $N_{\partial\Omega}=800$ on the boundary $\partial\Omega$. We run NGD-based methods for $1000$ iterations and BFGS for $1.2\times 10^{5}$ iterations.
Figure 10: Kovasnay Flow discretized via PINNs. Top row, from left to right: relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$. Bottom row, from left to right: relative $\mathrm{H}^{1}(\Omega)$ error on $p$ vs. iterations (first $1000$ iterations), and relative $\mathrm{H}^{1}(\Omega)$ error on $p$ vs. wall-clock time.
Figure 11: Kovasnay Flow discretized via PINNs and trained with NGD using SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations.
E.4 Source recovery problem for the Poisson equation in 3D
Given a penalization parameter $\eta>0$ and interior observations $u_{d}\in\mathrm{L}^{2}(\Omega)$, we consider the inverse problem of recovering both the solution $u$ and the source term $f$ from data by minimizing the Tikhonov-regularized objective:
$\inf_{(u,f)\in\mathrm{H}\times\mathrm{L}^{2}(\Omega)}J(u,f)=\frac{1}{2}\left\|{u-u_{d}}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{\eta^{2}}{2}\left\|{f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2},$ (33)
subject to the Poisson PDE constraint
$\begin{cases}-\Delta u=f&\text{in }\Omega=(0,1)^{3},\\ u=g_{D}&\text{on }\partial\Omega,\end{cases}$ (34)
where $\mathrm{H}=\overline{C^{\infty}(\overline{\Omega})}^{\left\|{\cdot}\right\|_{\mathrm{H}}}$ with $\left\|{u}\right\|_{\mathrm{H}}=\left\|{\Delta u}\right\|_{\mathrm{L}^{2}(\Omega)}+\left\|{u}\right\|_{\mathrm{L}^{2}(\partial\Omega)}$. See [72, Section 4.7] for well-posedness and theoretical background. While the solution $u$ in [72] yields $f=0$, in this paper we choose a more complex problem, where $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$ and the resulting source term to be reconstructed is not identically zero.
We discretize problem (3334) in strong form using a PINN formulation, where observations $u_{d}\in\mathrm{L}^{2}(\Omega)$ are in practice replaced by $N_{\mathrm{data}}$ point measurements. The function spaces are $\mathrm{V}=\mathrm{H}\times\mathrm{L}^{2}(\Omega)$, $\mathrm{W}=\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\partial\Omega)\mathrm{L}^{2}(\Omega)$, and the linear operator
$\mathcal{A}u=\left(\Delta u+f,u,\left.{u}\right|_{\partial\Omega},\eta f\right).$
The corresponding least squares loss is
$E(u,f)=\frac{1}{2}\left\|{\Delta u+f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(\partial\Omega)}^{2}+\frac{1}{2}\left\|{u-u_{d}}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{\eta^{2}}{2}\left\|{f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2},$ (35)
with induced metric
${g}\left({(u,f)},\,{(v,h)}\right)=\left({\Delta u+f},\,{\Delta v+h}\right)_{\mathrm{L}^{2}(\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\partial\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\Omega)}+\eta^{2}\left({f},\,{h}\right)_{\mathrm{L}^{2}(\Omega)}.$ (36)
We take $(u,f)=(u_{\boldsymbol{\theta}_{1}},f_{\boldsymbol{\theta}_{2}})$ two neural networks with weights combined into a single vector $\boldsymbol{\theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}]$, and approximate $L(\boldsymbol{\theta})=E(u_{{\boldsymbol{\theta}}_{1}},f_{{\boldsymbol{\theta}}_{2}})$ using Monte Carlo quadrature and point observations of $u_{d}$. Although $\eta>0$ is theoretically required to ensure the minimum of (35) exists in $\mathrm{V}$, we set $\eta=0$ in our numerical experiments.
Computational details
We employ fully connected neural networks $u_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$ with $\tanh$ activation and two hidden layers of width $32$. The term $\frac{1}{2}\left\|{u-u_{d}}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}$ in the energy is discretized using $N_{\mathrm{data}}=40$ noiseless observations of the true solution at random points in $\Omega$. The remaining integrals in the loss and metric are discretized via Monte Carlo quadrature with $N_{\Omega}=3000$ points uniformly sampled in the interior $\Omega$ and $N_{\partial\Omega}=800$ on the boundary $\partial\Omega$. We run NGD-based methods for $500$ iterations and BFGS for $1.2\times 10^{5}$ iterations.
Figure 12: Source recovery problem for the Poisson equation in 3D discretized via PINNs. Top row, from left to right: relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations (first $500$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$. Bottom row, from left to right: relative $\mathrm{L}^{2}(\Omega)$ error on $f$ vs. iterations (first $500$ iterations), and relative $\mathrm{L}^{2}(\Omega)$ error on $f$ vs. wall-clock time.
Figure 13: Source recovery problem for the Poisson equation in 3D discretized via PINNs and trained with NGD using SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations.
Appendix F Additional Numerical Experiments: Variational Formulations of PDEs
We now consider variational formulations of PDEs, where the solution is approximated by a neural network $u_{\boldsymbol{\theta}}$. We consider formulations closely related to Variational PINNs (VPINNs) [41, 40] and the Deep Ritz method [25]. In all examples, we interpolate the neural network onto a conforming finite element (FE) space, following [7]. This approach offers several advantages. First, Dirichlet boundary conditions can be approximately enforced by applying them to the FE interpolant using standard FE techniques, avoiding the need to impose them directly on the neural network output [10, 9]. Second, it alleviates integration challenges, a known issue in neural network-based PDE solvers [61], especially in the Deep Ritz method [64]. Finally, spatial derivatives can be computed at the FE level, eliminating the need for automatic differentiation with respect to the input of the neural network. Further details are provided in [7, 6].
In all experiments involving variational formulations, we set the parameters in Algorithm 1 as follows: $\ell_{0}=10$, $\ell_{\max}=500$, $\gamma=30$, $\mathtt{maxit}=20$, and $\kappa=0.1$. The neural networks are interpolated on a finite element space of degree $d=4$ on a uniform quadrilateral mesh with mesh size $h=1/30$, yielding approximately $14,500$ degrees of freedom. Integrals in the loss and in the metric are evaluated by interpolation onto the FE space and Gauss-Jacobi quadrature of order $82d=$ 8. Errors are computed with a Gauss-Jacobi quadrature rule of order $164d=$ 16. We rely on the Julia packages Lux.jl [57] for neural networks, and Optim.jl [52] for optimization.
F.1 Poisson problem in 2D – Deep Ritz formulation
We consider the Poisson problem on the 2D unit square
$\begin{cases}-\Delta u=f&\Omega=(0,1)^{2},\\ u=g_{D}&\partial\Omega,\end{cases}$ (37)
where $f\in\mathrm{H}^{-1}(\Omega)$ and $g_{D}\in\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$. As exact solution we take $u(x,y)=\sin(\pi x)\sin(\pi y)$.
All equalities in (37) are understood in the variational sense. That is, given a lifting $\tilde{g}_{D}\in\mathrm{H}^{1}(\Omega)$ such that $\left.{\tilde{g}_{D}}\right|_{\partial\Omega}=g_{D}$, we seek $u_{0}\in\mathrm{H}^{1}_{0}(\Omega)$ satisfying
$a(u_{0},v)=\int_{\Omega}fv\,\mathrm{d}\mathbf{x}-a(\tilde{g}_{D},v)\quad\forall v\in\mathrm{H}^{1}_{0}(\Omega),\qquad\text{where}\quad a(u,v)=\int_{\Omega}\nabla u\cdot\nabla v\,\mathrm{d}\mathbf{x}.$ (38)
The solution is then obtained as $u=u_{0}+\tilde{g}_{D}$. The spaces for the discretization are therefore $\mathrm{V}=\mathrm{W}=\mathrm{H}^{1}_{0}(\Omega)$, and the linear operator is given by
$\mathcal{A}:\mathrm{H}^{1}_{0}(\Omega)\to\mathrm{H}^{-1}(\Omega),\qquad\left\langle{\mathcal{A}u},\,{v}\right\rangle_{\mathrm{H}^{-1}\times\mathrm{H}^{1}_{0}}=a(u,v)=\int_{\Omega}\nabla u\cdot\nabla v\,\mathrm{d}\mathbf{x}.$
Since $a$ is a symmetric and coercive bilinear form on $\mathrm{V}\times\mathrm{V}$, we can employ the Deep Ritz formulation and consider the minimization problem
$\begin{split}\min_{u_{0}\in\mathrm{H}^{1}_{0}(\Omega)}E(u_{0})&\coloneqq\int_{\Omega}\frac{1}{2}\left\|{\nabla u_{0}}\right\|^{2}-fu_{0}+\nabla\tilde{g}_{D}\cdot\nabla u_{0}\,\mathrm{d}\mathbf{x},\end{split}$ (39)
with corresponding metric is $g=a$.
To discretize (37), we approximate the solution $u=u_{0}+\tilde{g}_{D}$ as $u\approx\Pi u_{\boldsymbol{\theta}}+\tilde{g}_{D}$, where $u_{\boldsymbol{\theta}}$ is a neural network, $\Pi$ is the interpolation operator onto the non-Dirichlet degrees of freedom of the FE space, and an approximate lifting $\tilde{g}_{D}$ is obtained by interpolating $g_{D}$ on the boundary degrees of freedom using standard FE procedures. We then minimize the loss $L(\boldsymbol{\theta})=E(\Pi u_{\boldsymbol{\theta}})$.
Computational details
We use a fully connected neural network $u_{\boldsymbol{\theta}}$ with $\tanh$ activation and two hidden layers of width $32$. We run NGD-based methods for $500$ iterations and BFGS for $1.2\times 10^{5}$ iterations. Since the loss can become negative, for NGD and NGD with CG we cannot employ the heuristic $\mu=\min(L(\boldsymbol{\theta}),10^{-5})$ as suggested in [72], and therefore we set $\mu=10^{-9}$. In NyströmNGD we take $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-9})$.
Results
On this test case, NyströmNGD and BFGS achieve comparable final accuracy, as shown in Figure 14, but NyströmNGD requires a fraction of the iterations and is approximately 10×10$\times$10 × faster in wall-clock time. This efficiency arises from the excellent low-rank approximability of the Gramian, evident in Figure 15, which leads the Nyström rank parameter $\ell$ to stabilize around $80$, far below the parametric dimension $p$. NGD and NGD with CG are not competitive in either iteration count or runtime, and both plateau at lower accuracy.
Figure 14: Poisson equation in 2D discretized via Deep Ritz and finite element interpolation. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 15: Poisson equation in 2D discretized via Deep Ritz and finite element interpolation, trained with NGD using an SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
F.2 Diffusion-reaction-transport problem in 2D
We consider the diffusion-reaction-transport problem
$\begin{dcases}-\nabla\cdot(\kappa\nabla u)+\boldsymbol{\beta}\cdot\nabla u+\sigma u=F&\Omega=(0,1)^{2},\\ u=g&\Gamma_{D}\subset\partial\Omega,\\ \kappa\nabla u\cdot\mathbf{n}=\psi&\Gamma_{N}\subset\partial\Omega,\end{dcases}$ (40)
where $\kappa(x,y)=2+\sin(x+2y)$, $\boldsymbol{\beta}(x,y)=\begin{bmatrix}\sqrt{x-y^{2}+5}&\sqrt{y-x^{2}+5}\end{bmatrix}^{\top}$, and $\sigma(x,y)=e^{\frac{x}{2}-\frac{y}{3}}+2$. The source term $F$ is chosen so that the exact solution is
$u(x,y)=\sin(3.2x(x-y))\cos(4.3y+x)+\sin(4.6(x+2y))\cos(2.6(y-2x)),$
as in [10, 7]. The Dirichlet boundary $\Gamma_{D}$ consists of the left and right sides of the domain, while the Neumann boundary $\Gamma_{N}$ corresponds to the top and bottom edges.
The function spaces are $\mathrm{V}=\mathrm{W}=\mathrm{H}^{1}_{0,\Gamma_{D}}(\Omega)$, and the variational formulation of (40) is defined as $\mathcal{A}u=f$ where
$\begin{split}\left\langle{\mathcal{A}u},\,{v}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}}&=a(u,v)=\int_{\Omega}(\kappa\nabla u)\cdot\nabla v+(\boldsymbol{\beta}\cdot\nabla u)\,v+\sigma\,u\,v\,\,\mathrm{d}\mathbf{x}\\ \left\langle{f},\,{v}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}}&=\int_{\Omega}F\,v\,\,\mathrm{d}\mathbf{x}+\int_{\Gamma_{N}}\psi\,v\,\,\mathrm{d}\mathbf{s}\end{split}$
Since $\mathcal{A}$ is in general not symmetric positive definite, we adopt the least-squares formulation
$\min_{u\in\mathrm{H}^{1}_{0}(\Omega)}E(u)\coloneqq\frac{1}{2}\left\|{\mathcal{A}u-f}\right\|_{\mathrm{H}^{-1}_{\Gamma_{D}}}^{2}=\left\langle{\mathcal{A}u-f},\,{\mathcal{R}_{\mathrm{V}}^{-1}(\mathcal{A}u-f)}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}},$ (41)
with associated metric ${g}\left({v},\,{w}\right)\coloneqq\left\langle{\mathcal{A}w},\,{\mathcal{R}_{\mathrm{V}}^{-1}\mathcal{A}v}\right\rangle$. Here, $\mathcal{R}_{\mathrm{V}}^{-1}=(-\Delta)^{-1}:\mathrm{H}^{-1}_{\Gamma_{D}}\to\mathrm{H}^{1}_{0,\Gamma_{D}}$ denotes the inverse Riesz operator, which involves solving a Poisson equation.
To discretize (41) and the corresponding metric, we interpolate the neural network $u_{\boldsymbol{\theta}}$ onto a conforming finite element space $\mathrm{V}_{h}\subset\mathrm{V}$, following [7], and minimize $L(\boldsymbol{\theta})=E(u_{\boldsymbol{\theta}})$. Applying the discrete Riesz operator $\mathcal{R}_{\mathrm{V}_{h}}^{-1}$ corresponds to solving a Poisson problem with FEM. To reduce computational cost, one could instead use a spectrally equivalent approximation, such as a few cycles of a geometric multigrid preconditioner. However, for simplicity, we solve the Poisson problem directly using FEM, computing a Cholesky factorization of the Laplacian matrix once.
Computational details
We use a fully connected neural network $u_{\boldsymbol{\theta}}$ with $\tanh$ activation and two hidden layers of width $32$. We run NGD-based methods for $1000$ iterations and BFGS for $1.2\times 10^{5}$ iterations. For NGD and NGD with CG we use $\mu=\min(10^{-5},L(\boldsymbol{\theta}))$ as suggested in [72]. For NyströmNGD, to enforce stronger regularization in early iterations, we take $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-8}\cdot L(\boldsymbol{\theta})^{2})$.
Results
As illustrated in Figure 16, NyströmNGD reaches convergence in few iterations and achieves better accuracy compared to other NGD methods. The gap in convergence speed relative to BFGS is substantial in terms of both iterations and time. The smaller difference in wall-clock time between NyströmNGD and BFGS, compared to the previous variational example, can be attributed to the higher Nyström rank parameter $\ell$ at which NyströmNGD stabilizes. Although $\ell$ remains much lower than the parametric dimension $p$, it is higher than in earlier examples. This is due to the slower, yet still strong, decay of the singular values of the Gramian, as shown in Figure 17.
Figure 16: Diffusion-reaction-transport equation in 2D discretized via Finite Element Interpolated Neural Networks. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 17: Diffusion-reaction-transport equation in 2D discretized via Finite Element Interpolated Neural Networks, trained with NGD using an SVD-based pseudoinverse. On the left, the decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. On the right, the relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims in the abstract and introduction are supported by experimental results, and the paper’s scope is clearly defined.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discusses its limitations primarily in the discussion section (Section 6) and also touches on them during the analysis of numerical experiments.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The paper provides a full set of assumptions for 3.2, and appropriately references assumptions for all other minor theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper includes detailed pseudocode of the algorithm in Algorithm 1 and full experimental details in Appendix E and Appendix D, ensuring reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: No data is used. The code will be publicly released on GitHub shortly before the conference, contingent on acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All necessary training and test details are provided in Section 5 and Appendix E.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The results of all numerical experiments are averaged over ten runs, with the median and the area between the first and third quartiles reported.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Computer resources, including execution time and type of compute workers, are reported in Section 5 for all numerical experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We guarantee that the research conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The paper focuses on a novel algorithm for training neural networks in Scientific Machine Learning, with no societal impacts to disclose.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No such risks are associated with this paper.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification: The paper does not use existing datasets. All packages and code are properly mentioned, and the corresponding papers are adequately credited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: No new assets are introduced in the paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.