### v3

# Accelerating Natural Gradient Descent for PINNs with Randomized Nyström Preconditioning

Abstract
Natural Gradient Descent (NGD) has emerged as a promising optimization algorithm for training neural network-based solvers for partial differential equations (PDEs), such as Physics-Informed Neural Networks (PINNs). However, its practical use is often limited by the high computational cost of solving linear systems involving the Gramian matrix. While matrix-free NGD methods based on the conjugate gradient (CG) method avoid explicit matrix inversion, the ill-conditioning of the Gramian significantly slows the convergence of CG. In this work, we extend matrix-free NGD to broader classes of problems than previously considered and propose the use of Randomized Nyström preconditioning to accelerate convergence of the inner CG solver. The resulting algorithm demonstrates substantial performance improvements over existing NGD-based methods and other state-of-the-art optimizers on a range of PDE problems discretized using neural networks.
Key words: Natural Gradient, Nyström Preconditioning, Neural Networks, PDEs
1 Introduction
Neural network (NN) based approaches for solving partial differential equations (PDEs) have recently attracted significant attention as promising alternatives to traditional numerical methods for both forward and inverse problems. A common strategy is to incorporate the governing PDE directly into the loss function, as in Physics-Informed Neural Networks (PINNs) [29, 70, 75], Variational PINNs [49, 50], the Deep Ritz method [30], and various subsequent extensions [8, 38, 9, 25, 73]. These methods hold the promise of providing mesh-free approximations of PDE solutions and of exploiting the approximation power of neural networks [28, 23], particularly in high dimensions [18], positioning them as promising tools for high-dimensional, inverse, or complex domain problems. Despite this theoretical promise, the practical adoption and reliability of NN-based PDE solvers are often hindered by significant optimization challenges. Standard first-order algorithms, notably Adam [51], have been reported to often perform poorly, stagnating at solutions with low accuracy [55, 80, 71, 81]. This issue is generally attributed not to the expressiveness of neural networks, but to the ill-conditioning of the loss landscape [71, 24, 55]. Consequently, specialized optimization strategies tailored to PDE-related loss functions are increasingly being advocated [21, 65, 74].
To counteract these training pathologies, various strategies have been proposed. A body of work focuses on loss weighting schemes or adaptive selection of collocation points to balance contributions from different terms in the loss and reduce quadrature errors [80, 14, 79, 82, 84]. While these techniques can mitigate some failure modes, they do not fully resolve the underlying ill-conditioning introduced by discretizing a PDE operator. Only recently, more principled second-order optimization algorithms have begun to emerge [64, 65, 24, 86, 22, 71, 39, 46], often derived from preconditioning the underlying infinite-dimensional PDE operator [24, 65, 64, 39]. Among them, Natural Gradient Descent (NGD) has emerged as a particularly promising and unifying framework, as several earlier approaches can be interpreted as NGD with different choices of the optimization metric [66, 65].
NGD has demonstrated remarkable performance, achieving high accuracy in few iterations by using a search direction of the form $\mathbf{d}=-(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})^{-1}\nabla L(\boldsymbol{\theta})$, where $\boldsymbol{\theta}\in\mathbb{R}^{p}$ denotes the neural network parameters, $L$ is the loss function, $\mathbf{G}(\boldsymbol{\theta})$ the positive semidefinite Gramian matrix (see Section˜3.1), and $\mu>0$ a regularization parameter. However, solving linear systems with $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$ via direct methods incurs a cubic cost $\mathcal{O}(p^{3})$ in the number of network parameters, making NGD impractical for large-scale networks. To alleviate this, matrix-free strategies have been proposed [46, 86], where the system is solved approximately using the conjugate gradient (CG) method using only matrix-vector products (matvecs) with $\mathbf{G}(\boldsymbol{\theta})$, without assembling the matrix. In practice, however, $\mathbf{G}(\boldsymbol{\theta})$ is often ill-conditioned, as shown in Figure˜1, leading to slow CG convergence and motivating the need for preconditioning. In this work, we address this issue using tools from Randomized Numerical Linear Algebra (RandNLA) [60, 63].
The core contributions of this work are as follows. First, we provide a general methodology to efficiently compute matrix-vector products with the Gramian using automatic differentiation, extending beyond the specific metrics considered in prior works [65, 46, 86]. Second, we propose the use of randomized Nyström preconditioning [31] to construct a preconditioner for $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$. Derived from a low-rank approximation of the Gramian, this preconditioner leverages the empirically observed strong spectral decay, turning it from a challenge into an asset. Third, we introduce a novel optimization algorithm for neural network-based PDE solvers, termed NyströmNGD, which combines matrix-free NGD with Randomized Nyström preconditioning. Finally, we present extensive numerical experiments demonstrating significant acceleration and improved final accuracy compared to existing NGD-based methods and other state-of-the-art optimizers.
Matrix-free NGD implementations have already been explored for specific metrics in [46, 86], but none of these works address the ill-conditioning of the Gramian, which hampers CG convergence in practice. To address this bottleneck, our work bridges NGD with Randomized Numerical Linear Algebra (RandNLA), whose use in Machine Learning is well-established [32, 27, 78]. Randomized Nyström preconditioning [31] has been applied to PINNs in [71], but only as a final stage after Adam and L-BFGS, and applying the Nyström approximation, designed for symmetric positive semidefinite matrices, to the Hessian, which may be indefinite. RandNLA has been applied to NGD in the contemporaneous and independent work [61], but there RandNLA is used in the so-called sketch-and-solve paradigm, whereas we employ it to construct a preconditioner. Moreover, in that work the Nyström approximation is not employed even if the Gramian is symmetric positive semidefinite, and the approach is not fully matrix-free. Alternative notions of preconditioning, where the preconditioner is embedded into the loss function, are explored in [56, 8]. Kronecker-Factored Approximate Curvature (K-FAC) [22] has also been used to approximate the Gramian [22], but not as a preconditioner, and it requires problem-specific adaptations.
Paper outline
The remainder of the paper is structured as follows. In Section˜2, we introduce the abstract formulation of the problem and the neural network framework. Section˜3 reviews Natural Gradient Descent and presents its matrix-free formulation. In Section˜4, we provide the necessary background on the randomized Nyström approximation, and in Section˜4.2 we present our proposed algorithm, NyströmNGD, in full detail. Section˜5 contains numerical experiments on a range of PDE-related problems, comparing our method against state-of-the-art optimizers. Finally, Section˜6 summarizes our findings and outlines directions for future research.
Figure 1: Spectral decay of the Gramian matrix when training a PINN for the 3D Poisson problem with exact solution $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$, using Energy Natural Gradient Descent [64]. Left: spectral decay of the Gramian $\mathbf{G}(\boldsymbol{\theta})$ at different iterations, with each color corresponding to a different iteration. Right: $\mathrm{H}^{1}(\Omega)$ relative error during training.
2 Preliminaries
2.1 Notation
Hilbert and Banach spaces are indicated using roman capital letters. Linear operators are denoted with calligraphic letters. Vectors in $\mathbb{R}^{n}$ are in boldface, and matrices use boldface capital letters.
Let $\mathrm{V}$ be a real Hilbert space. Its inner product and norm are denoted by $\left({\cdot},\,{\cdot}\right)_{\mathrm{V}}$ and $\left\|{\cdot}\right\|_{\mathrm{V}}$. Its (topological) dual is indicated $\mathrm{V}^{\prime}$, and the duality pairing by $\left\langle{\cdot},\,{\cdot}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}$. The space of bounded linear operators from $\mathrm{V}$ to another Hilbert space $\mathrm{W}$ is denoted by $\mathcal{L}(\mathrm{V},\mathrm{W})$. We say that $\mathcal{A}\in\mathcal{L}(\mathrm{V},\mathrm{V}^{\prime})$ is symmetric positive semidefinite (SPSD) if the associated bilinear form $a(u,v)=\left\langle{\mathcal{A}u},\,{v}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}$ is symmetric and satisfies $a(u,u)\geq 0$ for all $u\in\mathrm{V}$. It is symmetric positive definite (SPD) if $a$ is also coercive, i.e., if $\exists\,\alpha>0$ such that $a(u,u)\geq\alpha\left\|{u}\right\|_{\mathrm{V}}^{2}$ for all $u\in\mathrm{V}$.
Let $\left(\mathcal{M},g\right)$ be a smooth Riemannian manifold. The tangent space at $u\in\mathcal{M}$ is denoted by $\mathrm{T}_{u}\mathcal{M}$, with inner product and norm given by ${g_{u}}\left({\cdot},\,{\cdot}\right)$ and $\left\|{\cdot}\right\|_{g_{u}}$, respectively. For a smooth cost function $E:\mathcal{M}\to\mathbb{R}$, the Riemannian gradient and Hessian are denoted by $\mathrm{grad}^{g}E$ and $\mathrm{Hess}^{g}E$, respectively. The Euclidean gradient of a function $E:\mathbb{R}^{n}\to\mathbb{R}$ is denoted by $\nabla E$.
2.2 Abstract Setup
We consider a general minimization problem of the form
$\inf_{u\in\mathrm{V}}E(u),$ (1)
where $\mathrm{V}$ is a Hilbert space endowed with a possibly point-dependent metric $g$ which gives to $\mathrm{V}$ the structure of a Riemannian manifold, and $E:\mathrm{V}\to\mathbb{R}$ is a smooth function, referred to as the energy.
To approximate its solution, we introduce a parametrized ansatz
$\mathcal{N}=\left\{u_{\boldsymbol{\theta}}\coloneqq\rho(\boldsymbol{\theta})\,:\,\boldsymbol{\theta}\in\Theta\right\}\subseteq\mathrm{V},$
where $\rho:\Theta\to\mathrm{V}$ is a smooth map and $\Theta\subseteq\mathbb{R}^{p}$ is open. We then consider the finite-dimensional optimization problem
$\inf_{\boldsymbol{\theta}\in\Theta}L(\boldsymbol{\theta})\coloneqq\inf_{\boldsymbol{\theta}\in\Theta}E(u_{\boldsymbol{\theta}})=\inf_{\boldsymbol{\theta}\in\Theta}(E\circ\rho)(\boldsymbol{\theta}).$ (2)
Here, $u_{\boldsymbol{\theta}}$ could be a neural network with parameters $\boldsymbol{\theta}$, though the framework applies more generally. We will refer to $\Theta\subseteq\mathbb{R}^{p}$ as the parameter space, and to $\mathrm{V}$ as the function space. Note that while the original optimization problem (1) may be infinite dimensional, the ansatz is parametrized by a finite-dimensional $\boldsymbol{\theta}\in\mathbb{R}^{p}$.
2.2.1 Discretization of PDEs
The focus of this article is on problems arising from the discretization of PDEs. In this case, $\mathrm{V}$ is a Hilbert space of functions defined on a domain $\Omega\subset\mathbb{R}^{d}$, e.g., $\mathrm{V}=\mathrm{H}^{1}(\Omega)$, and $E$ is derived from the strong or weak formulation of a PDE on $\Omega$. This abstract setting (2) encompasses, for instance, the Deep Ritz method [30], PINNs [70, 29], Variational PINNs (VPINNs)[49], and other PINNs variants [50, 8, 38, 25, 73].
As a general framework for PDEs we consider the following. Let $\mathrm{V}$ be Hilbert space and let $\mathrm{X},\mathrm{Y}$ be Banach spaces. We seek $u\in\mathrm{V}$ such that
$\begin{cases}\mathcal{F}(u)=f,&\text{in }\Omega,\\ \mathcal{B}(u)=g,&\text{on }\partial\Omega,\end{cases}$ (3)
where $\mathcal{F}:\mathrm{V}\to\mathrm{X}$ is a differential operator, $f\in\mathrm{X}$ is the source term, $\mathcal{B}:\mathrm{V}\to\mathrm{Y}$ is a boundary operator, and $g\in\mathrm{Y}$ is the boundary data. The energy can then be defined as
$E(u)=\frac{1}{2}\left\|{\mathcal{F}(u)-f}\right\|_{\mathrm{X}}^{2}+\frac{1}{2}\left\|{\mathcal{B}(u)-g}\right\|_{\mathrm{Y}}^{2},$
where possibly weighting factors can be included, and data misfit terms can be added.
The choice of function spaces and well-posedness depend on the specific PDE and on whether a strong or weak formulation is used. For example, consider the Poisson equation with homogeneous Dirichlet boundary conditions:
$\begin{cases}-\Delta u=f,&\text{in }\Omega,\\ u=g,&\text{on }\partial\Omega.\end{cases}$ (4)
For a strong formulation, once can employ the spaces $\mathrm{V}=\mathrm{H}^{2}(\Omega)$, $\mathrm{X}=\mathrm{L}^{2}(\Omega)$, and $\mathrm{Y}=\mathrm{L}^{2}(\partial\Omega)$, with energy
$E(u)=\frac{1}{2}\left\|{-\Delta u-f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{1}{2}\left\|{u-g}\right\|_{\mathrm{L}^{2}(\partial\Omega)}^{2}.$
If instead a weak formulation is used, the natural spaces are $\mathrm{V}=\mathrm{H}^{1}_{0}(\Omega)$, $\mathrm{X}=\mathrm{V}^{\prime}=\mathrm{H}^{-1}(\Omega)$, and $\mathrm{Y}=\mathrm{H}^{1/2}(\partial\Omega)$. Given a lifting $\tilde{g}\in\mathrm{H}^{1}(\Omega)$ of the Dirichlet data, the energy is
$E(u)=\frac{1}{2}\left\|{\mathcal{A}u-\tilde{f}}\right\|_{\mathrm{H}^{-1}(\Omega)}^{2},\qquad\text{where}\qquad\left\langle{\mathcal{A}u},\,{v}\right\rangle_{\mathrm{H}^{-1}\times\mathrm{H}^{1}_{0}}=\int_{\Omega}\nabla u\cdot\nabla v,\quad\tilde{f}=f-\mathcal{A}\tilde{g}$.
2.3 Neural Networks
Neural networks (NNs) serve as the primary parametrized ansatz in this work. For simplicity, we focus on feedforward networks, though our methods are architecture-agnostic and apply to other network types as well. A neural network with input dimension $d$ and output dimension $d^{\prime}$ is a function $u_{\boldsymbol{\theta}}^{\mathcal{NN}}:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ defined as the composition affine and nonlinear transformations. Let $L$ be the number of hidden layers, let $n_{0}=d,n_{L+1}=d^{\prime}$, and $n_{1},\dots,n_{L}$ be the number of neurons in each hidden layer, and let $\sigma$ be a nonlinear activation function. The NN is defined as:
$\begin{split}\mathbf{x}_{0}&=\mathbf{x}\in\mathbb{R}^{d},\\ \mathbf{x}_{\ell}&=\sigma(\mathbf{W}_{\ell}\mathbf{x}_{\ell-1}+\mathbf{b}_{\ell})\in\mathbb{R}^{n_{\ell}},\hskip 18.49988pt\text{for }\ell=1,\dots,L,\\ u_{\boldsymbol{\theta}}^{\mathcal{NN}}(\mathbf{x})=\mathbf{x}_{L}&=\mathbf{W}_{L+1}\mathbf{x}_{L}+\mathbf{b}_{L+1}\in\mathbb{R}^{d^{\prime}},\end{split}$
where $\mathbf{W}_{\ell}\in\mathbb{R}^{n_{\ell}\times n_{\ell-1}}$ and $\mathbf{b}_{\ell}\in\mathbb{R}^{n_{\ell}}$ are the weights and biases of layer $\ell$, respectively. The parameters of the network are $\boldsymbol{\theta}=[\mathbf{W}_{1},\mathbf{b}_{1},\dots,\mathbf{W}_{L+1},\mathbf{b}_{L+1}]\in\mathbb{R}^{p}$, stacked into a single vector, with $p=\sum_{\ell=1}^{L+1}(n_{\ell}n_{\ell-1}+n_{\ell})$. No activation is applied in the last layer.
2.3.1 Finite Element Interpolated Neural Networks (FEINNs)
While standard NNs are typically used with collocation-based methods like PINNs, to deal with variational formulations of PDEs, we consider Finite Element Interpolated Neural Networks (FEINNs) [8, 7, 10], where a neural network is interpolated onto a finite element (FE) space.
Let $\Omega\subset\mathbb{R}^{d}$ be a Lipschitz polyhedral domain with a conforming simplicial mesh $\mathcal{T}_{h}$, and let $\mathrm{V}_{h}\subset\mathrm{V}=\mathrm{H}^{1}(\Omega)$ be the FE space with $\mathrm{V}_{h,0}=\mathrm{V}_{h}\cap\mathrm{H}^{1}_{0}(\Omega)$. Denote the Lagrangian basis by $\{\varphi_{i}\}_{i=1}^{N}$, where $\varphi_{i}$ corresponds to node $\mathbf{x}_{i}$, and let $\Pi_{h}:C^{0}(\Omega)\to\mathrm{V}_{h,0}$ be the FE interpolation operator obtained by evaluation at interior nodes. Given a discrete lifting $\bar{u}_{h}$ of Dirichlet data, the FEINN is defined as
$u_{\boldsymbol{\theta}}=\bar{u}_{h}+\Pi_{h}u_{\boldsymbol{\theta}}^{\mathcal{NN}}=\bar{u}_{h}n+\sum_{i\,:\,\mathbf{x}_{i}\in\mathring{\Omega}}u_{\boldsymbol{\theta}}^{\mathcal{NN}}(\mathbf{x}_{i})\varphi_{i}.$
For simplicity, we assumed that Dirichlet boundary conditions are imposed on the whole boundary, but mixed boundary conditions can also be accommodated in the framework. Further details are provided in [8, 7].
This approach offers several advantages. First, Dirichlet boundary conditions can be approximately enforced by applying them to the FE interpolant using standard FE techniques, avoiding the need to impose them directly on the neural network output [10, 11]. Second, it alleviates integration challenges, a known issue in neural network-based PDE solvers [72], especially in the Deep Ritz method [76]. Finally, spatial derivatives can be computed at the FE level, eliminating the need for automatic differentiation with respect to the input of the neural network.
3 Matrix-free Natural Gradient Descent
3.1 Natural Gradient Descent
The Natural Gradient Descent (NGD) method [3, 59, 66, 65] targets optimization problems of the form (2) following the first-optimize-then-discretize paradigm. NGD moves along the steepest descent direction according to the metric on $\mathrm{V}$, the “natural” one for the function $u_{\boldsymbol{\theta}}$, and pulls it back to the parameter space.
Let $\boldsymbol{\theta}$ be the current iterate. The steepest descent direction in function space is $d=-\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}})$, the Riemannian gradient of $E$ in the metric $g$, and the corresponding update
$u\leftarrow u_{\boldsymbol{\theta}}-\alpha\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}}).$
In NGD, the parameter update direction is $\mathbf{d}$, where the natural gradient $\mathbf{d}\in\mathbb{R}^{p}$ is such that the first-order expansion
$u_{\boldsymbol{\theta}-\alpha\mathbf{d}}\approx u_{\boldsymbol{\theta}}-\alpha\mathrm{D}\rho(\boldsymbol{\theta})[\mathbf{d}]=u_{\boldsymbol{\theta}}-\alpha\sum_{j=1}^{p}d_{j}\cdot\partial_{\theta_{j}}u_{\boldsymbol{\theta}},$
best approximates (in the metric $g$) the update in function space, i.e.,
$\mathbf{d}=\operatorname*{arg\,min}_{\mathbf{w}\in\mathbb{R}^{p}}\left\|{\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}})-\mathrm{D}\rho(\boldsymbol{\theta})[\mathbf{w}]}\right\|_{g_{u_{\boldsymbol{\theta}}}}.$
This corresponds to the $g_{u_{\boldsymbol{\theta}}}$-orthogonal projection of $\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}})$ onto the model’s generalized tangent space
$\mathrm{T}_{\boldsymbol{\theta}}\mathcal{N}=\mathrm{Span}\left(\left\{\partial_{\theta_{1}}u_{\boldsymbol{\theta}},\dots,\partial_{\theta_{p}}u_{\boldsymbol{\theta}}\right\}\right)\subset\mathrm{V}.$
If $\partial_{\theta_{1}}u_{\boldsymbol{\theta}},\dots,\partial_{\theta_{p}}u_{\boldsymbol{\theta}}$ are not linearly independent, the minimum is not unique, and $\mathbf{d}$ of minimal norm is taken. As shown, e.g, in [3, 66], the natural gradient can be computed as
$\mathbf{d}=\mathbf{G}(\boldsymbol{\theta})^{\dagger}\nabla_{\boldsymbol{\theta}}L(\boldsymbol{\theta}),\qquad\left[\mathbf{G}(\boldsymbol{\theta})\right]_{ij}={g_{u_{\boldsymbol{\theta}}}}\left({\partial_{\theta_{i}}u_{\boldsymbol{\theta}}},\,{\partial_{\theta_{j}}u_{\boldsymbol{\theta}}}\right),$ (5)
where $\mathbf{G}(\boldsymbol{\theta})\in\mathbb{R}^{p\times p}$ is often referred to as the Gramian [85, 65] or Fisher information matrix [3, 59], and $\mathbf{G}(\boldsymbol{\theta})^{\dagger}$ denotes its Moore–Penrose pseudoinverse. As observed in the literature [85, 13], it is often beneficial to add a damping term to improve numerical stability and convergence, yielding the regularized direction
$\mathbf{d}=\left(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I}\right)^{-1}\nabla_{\boldsymbol{\theta}}L(\boldsymbol{\theta}),$ (6)
where $\mu>0$ is a regularization parameter, typically adapted during the optimization process. Once the (regularized) natural gradient descent is computed, the parameters are updated as
$\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\alpha\mathbf{d}=\boldsymbol{\theta}-\alpha(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})^{-1}\nabla_{\boldsymbol{\theta}}L(\boldsymbol{\theta}),$
where the step-size $\alpha$ is typically chosen via a line-search procedure.
3.2 A matrix-free formulation
Leaving aside the resolution of the least squares problem, even assembling the Gramian $\mathbf{G}(\boldsymbol{\theta})$ in (5) can be computationally prohibitive in both flops and memory. Indeed, $\mathbf{G}(\boldsymbol{\theta})$ is a $p\times p$ matrix, where $p$ is the number of network parameters, which can be in the order tens of thousands even for relatively small neural networks. To mitigate this issue, we consider a matrix-free approach: $\mathbf{G}(\boldsymbol{\theta})$ is never explicitly assembled, and only matrix-vector products are computed as needed. This is particularly efficient when solving linear systems with $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$ using iterative methods such as Conjugate Gradient (CG), which require only matvecs.
Prior work has demonstrated that matrix-vector products with the Gramian (5) can be computed via automatic differentiation for specific metrics [46, 65, 86]. Here, we show that autodiff-based matvecs extend to broader metric classes. Specifically, we show that matrix-vector products with the Gramian (5) can be computed via automatic differentiation, provided the metric satisfies the following assumption and it is discretized via quadrature.
Assumption 3.1.
Let $\Omega\subset\mathbb{R}^{d}$ be a domain, and let $\mathrm{V}$ be a Hilbert space of functions on $\Omega$. Assume that the metric $g_{u}$ takes the form
${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\mathcal{F}_{u}v(\mathbf{x})\cdot\mathcal{F}_{u}w(\mathbf{x})\,\varphi_{u}(\mathbf{x})\,\mathrm{d}\mathbf{x},$ (7)
for all $v,w\in\mathrm{T}_{u}\mathrm{V}\subseteq\mathrm{V}$, where $\mathcal{F}_{u}:\mathrm{V}\to\mathrm{L}^{2}(\Omega;\varphi_{u})^{k}$ is linear and continuous.
Theorem 3.2.
Let ˜3.1 hold, and let $\mathtt{sg}:\mathbb{R}^{p}\to\mathbb{R}^{p}$ denote the stop-gradient operator, formally defined by $\mathtt{sg}(\boldsymbol{\theta})=\boldsymbol{\theta}$ and $\nabla\mathtt{sg}(\boldsymbol{\theta})=\mathbf{0}$. If the integral (7) is approximated using a quadrature rule with points $\{\mathbf{x}_{r}\}_{r=1}^{q}$ and weights $\{w_{r}\}_{r=1}^{q}$, then
[…]
where $\mathbf{J}_{\boldsymbol{\theta}}$ denotes the Jacobian with respect to $\boldsymbol{\theta}$.
Proof.
Let $\mathbf{e}_{h}\in\mathbb{R}^{p}$ denote the $h$-th vector of the canonical basis, i.e., the vector that has a $1$ in the $h$-th entry and 0 in all other entries. Using the linearity and continuity of $\mathcal{F}_{u_{\boldsymbol{\theta}}}$,
$\begin{split}\mathcal{F}_{u_{\boldsymbol{\theta}}}\partial_{\theta_{h}}u_{\boldsymbol{\theta}}&=\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}\partial_{\theta_{h}}u_{\boldsymbol{\theta}}=\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}^{k}\left(\lim_{t\to 0}\frac{u_{\boldsymbol{\theta}+t\mathbf{e}_{h}}-u_{\boldsymbol{\theta}}}{t}\right)\\ &=\lim_{t\to 0}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}\left(\frac{u_{\boldsymbol{\theta}+t\mathbf{e}_{h}}-u_{\boldsymbol{\theta}}}{t}\right)=\lim_{t\to 0}\frac{\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}+t\mathbf{e}_{h}}-\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}}{t}\\ &=\partial_{\theta_{h}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}.\end{split}$
Note that the use of the stop-gradient operator is crucial to ensure that we do not differentiate $\mathcal{F}_{u_{\boldsymbol{\theta}}}$ with respect to $\boldsymbol{\theta}$. Hence,
$\begin{split}[\mathbf{G}(\boldsymbol{\theta})]_{ij}&={g_{u_{\boldsymbol{\theta}}}}\left({\partial_{\theta_{i}}u_{\boldsymbol{\theta}}},\,{\partial_{\theta_{j}}u_{\boldsymbol{\theta}}}\right)=\int_{\Omega}\mathcal{F}_{u_{\boldsymbol{\theta}}}\partial_{\theta_{i}}u_{\boldsymbol{\theta}}\cdot\mathcal{F}_{u_{\boldsymbol{\theta}}}\partial_{\theta_{j}}u_{\boldsymbol{\theta}}(\mathbf{x})\,\varphi_{u_{\boldsymbol{\theta}}}\,\mathrm{d}\mathbf{x}\\ &=\int_{\Omega}\partial_{\theta_{i}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}\cdot\partial_{\theta_{j}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}\,\varphi_{u_{\boldsymbol{\theta}}}\,\mathrm{d}\mathbf{x},\end{split}$
or, in matrix form,
$\mathbf{G}(\boldsymbol{\theta})=\left(\mathrm{D}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}\right)^{*}\,\mathrm{D}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}},$
where the Hilbert adjoint $\left(\mathrm{D}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}\right)^{*}$ is taken with respect to the inner product on $\mathrm{L}^{2}(\Omega,\varphi_{u})^{k}$, and we identify a matrix with the induced linear map.
Applying quadrature with nodes $\{\mathbf{x}_{r}\}_{r=1}^{q}$ and weights $\{w_{r}\}_{r=1}^{q}$, we get
$\begin{split}[\mathbf{G}(\boldsymbol{\theta})]_{ij}&=\sum_{r=1}^{q}w_{r}\,\partial_{\theta_{i}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{r})\cdot\partial_{\theta_{j}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{r})\,\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{r})\end{split}$
and thus, in matrix form:
$\begin{split}\mathbf{G}(\boldsymbol{\theta})&=\sum_{r=1}^{q}w_{r}\,\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{r})\,\left(\mathbf{J}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{r})\right)^{\top}\mathbf{J}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{r})\\ &=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}(\boldsymbol{\theta})^{\top}\begin{bmatrix}w_{1}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{1})\mathbf{I}_{k}&&\\ &\ddots&\\ &&w_{q}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{q})\mathbf{I}_{k}\end{bmatrix}\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}(\boldsymbol{\theta}).\end{split}$
As a consequence of ˜3.2, matrix-vector products with $\mathbf{G}(\boldsymbol{\theta})$ can be computed using only products with the Jacobian and the transpose Jacobian of $\mathbf{F}$. These operations can be efficiently implemented via automatic differentiation, using forward-mode autodiff for Jacobian-vector products (JVPs) and backward-mode autodiff for vector-Jacobian products (VJPs). Note that most autodiff frameworks include a stop-gradient function.
Computational Cost
With automatic differentiation, the cost of each JVP and VJP is a small constant factor times the cost of evaluating $\mathbf{F}(\boldsymbol{\theta})$ [34, 58]. For a feedforward neural network, the evaluation cost scales as $\mathcal{O}(pq)$, where $p$ is the number of trainable parameters and $q$ is the number of quadrature points. Therefore, the cost of each matrix-vector product with $\mathbf{G}(\boldsymbol{\theta})$ is $\mathrm{C}_{\mathrm{mv}}=\mathcal{O}(pq)$. Note however that when executed on modern GPUs batch evaluation across $q$ points benefits from parallelization.
Remark 3.3.
More in general, it can be shown that if the metric $g_{u}$ takes the form
${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\mathcal{F}_{u}^{1}v(\mathbf{x})\cdot\mathcal{F}_{u}^{2}w(\mathbf{x})\,\varphi_{u}(\mathbf{x})\,\mathrm{d}\mathbf{x},$ (9)
where $\mathcal{F}_{u}^{i}:\mathrm{V}\to\mathrm{L}^{2}(\Omega;\varphi_{u})^{k}$ are linear and continuous for $i=1,2$, and the integral is approximated using a quadrature rule with points $\{\mathbf{x}_{r}\}_{r=1}^{q}$ and weights $\{w_{r}\}_{r=1}^{q}$, then
[…]
Remark 3.4.
˜3.1 encompasses standard metrics used in PDE-related function spaces. For instance, the $\mathrm{L}^{2}(\Omega)$ and $\mathrm{H}^{1}(\Omega)$ metrics correspond to the choices $\mathcal{F}_{u}v=v$ and $\mathcal{F}_{u}v=(\nabla v,v)$, respectively. It also includes the Gauss-Newton metric on $\mathrm{L}^{2}(\Omega)$, for which matrix-free NGD implementations have been proposed in [46, 65] for specific problems. However, ˜3.1 is significantly more general and covers a much broader class of metrics. In practice, it includes many of the metrics relevant in the context of PDEs, since these typically need to be expressible as integrals to be computationally tractable. For example, it includes the Newton metric ${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\nabla v\cdot\nabla w+3u^{2}vw$ in [65, Section 4.2], and the energy inner product for the Gross–Pitaevskii equation as considered in [41], ${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\nabla v\cdot\nabla w+Vvw+\beta\lvert u\rvert^{2}vw$.
3.2.1 A link with Operator Preconditioning and the case of Finite Elements Interpolated Neural Networks
The Gramian (5) can be interpreted in the context of operator preconditioning [44, 57, 52, 36]. Specifically, NGD can be seen as a preconditioned gradient descent method, with the Gramian serving as the preconditioner [24]. For a linear parametrized model $u_{\boldsymbol{\theta}}=\sum_{i=1}^{p}\theta_{i}\,\varphi_{i}$, the Gramian reduces to $\left[\mathbf{G}(\boldsymbol{\theta})\right]_{ij}={g_{u_{\boldsymbol{\theta}}}}\left({\varphi_{i}},\,{\varphi_{j}}\right)$. If $(\mathrm{V},\left({\cdot},\,{\cdot}\right)_{\mathrm{V}})$ is a Hilbert space and $g$ is induced by a positive definite linear operator $\mathcal{G}:\mathrm{V}\to\mathrm{V}^{\prime}$, then
$\mathbf{G}(\boldsymbol{\theta})]_{ij}=[\mathbf{G}]_{ij}=\left\langle{\mathcal{G}\varphi_{i}},\,{\varphi_{j}}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}},$
which is independent of $\boldsymbol{\theta}$. This coincides exactly with preconditioners arising from operator preconditioning in Galerkin discretizations of PDEs [44].
The connection with operator preconditioning highlights that the choice of metric crucially influences the convergence speed of NGD. In general, the link between NGD and Riemannian gradient descent [64] suggests selecting the metric $g$ to reduce the condition number of the Riemannian Hessian. Further details are provided in Appendix˜A, where we present a unified framework for metric selection.
In the case of FEINNs, ˜3.2 can be specialized to connect with operator preconditioning for finite elements. Assume, for simplicity, that the interior nodes of the mesh are the first $N_{I}$, so that the FEINN can be written as $u_{\boldsymbol{\theta}}=\bar{u}_{h}+\sum_{i=1}^{N_{I}}u_{\boldsymbol{\theta}}^{\mathcal{NN}}(\mathbf{x}_{i})\varphi_{i}$. Then the Gramian (5) takes the form
$\mathbf{G}(\boldsymbol{\theta})=\left(\mathbf{J}_{\boldsymbol{\theta}}\mathbf{P}(\boldsymbol{\theta})\right)^{\top}\mathbf{G}_{h}\left(\mathbf{J}_{\boldsymbol{\theta}}\mathbf{P}(\boldsymbol{\theta})\right),\qquad\text{where}\qquad\mathbf{P}(\boldsymbol{\theta})=\begin{bmatrix}u_{\boldsymbol{\theta}}^{\mathcal{NN}}(\mathbf{x}_{1})\\ \vdots\\ u_{\boldsymbol{\theta}}^{\mathcal{NN}}(\mathbf{x}_{N_{I}})\end{bmatrix}$,
where $\mathbf{G}_{h}\in\mathbb{R}^{N_{I}\times N_{I}}$ with $\mathbf{G}_{h}]_{ij}=\left\langle{\mathcal{G}\varphi_{i}},\,{\varphi_{j}}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}$ is the FEM operator preconditioner. Note that computing matvecs with the gramian requires only matvecs with $\mathbf{G}_{h}$, but not with its inverse, and JVPs and VJPs with $\mathbf{P}(\boldsymbol{\theta})$, which is an evaluation of the neural network at the interior nodes of the mesh.
4 NyströmNGD: Matrix-Free Natural Gradient Descent with Randomized Nyström Preconditioning
Given that matrix-vector products with the Gramian $\mathbf{G}=\mathbf{G}(\boldsymbol{\theta})$ can be computed efficiently, it is appealing to compute the regularized natural gradient (6) by solving the linear system with an iterative method. Since $\mathbf{G}+\mu\mathbf{I}$ is symmetric positive definite, the method of choice is Conjugate Gradient (CG). In practice, $\mathbf{G}$ is often ill-conditioned, causing slow convergence of CG unless a suitable preconditioner is employed. However, the matrix $\mathbf{G}$ lacks the sparsity or grid-based structure typical of FEM/FD matrices, which are crucial for classical preconditioners such as Incomplete Cholesky, multigrid/multilevel methods, and domain decomposition techniques. Instead, the main characteristics of $\mathbf{G}$ are that it is symmetric positive semidefinite, typically exhibits rapid singular value decay, and can only be accessed through matrix-vector products. In this setting, Randomized Nyström [77, 31] is, to the best of our knowledge, the most effective method currently available for constructing a preconditioner. In this section, we first review randomized Nyström preconditioning, then present our proposed method, NyströmNGD, which combines matrix-free NGD with randomized Nyström.
4.1 Randomized Nyström approximation and preconditioning
Let $\mathbf{G}\in\mathbb{R}^{p\times p}$ be symmetric positive semidefinite (PSD). The randomized Nyström approximation [60, Section 14] with a random test matrix $\boldsymbol{\Omega}\in\mathbb{R}^{p\times\ell}$ is
$\hat{\mathbf{G}}_{\mathrm{nys}}\coloneqq(\mathbf{G}\boldsymbol{\Omega})\left(\boldsymbol{\Omega}^{\top}\mathbf{G}\boldsymbol{\Omega}\right)^{\dagger}(\mathbf{G}\boldsymbol{\Omega})^{\top}.$ (11)
The randomness in $\boldsymbol{\Omega}$ ensures that $\hat{\mathbf{G}}_{\mathrm{nys}}$ approximates $\mathbf{G}$ well with [60, Section 14.4]. In this work, we take $\boldsymbol{\Omega}$ to be standard normal. Then the following guarantees hold.
Proposition 4.1.
[60, Theorem 14.1] Let $\mathbf{G}\in\mathbb{R}^{p\times p}$ be a symmetric positive semidefinite matrix with eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{p}\geq 0$. Let $\boldsymbol{\Omega}\in\mathbb{R}^{p\times\ell}$ be a standard Gaussian matrix with $\ell<p$. Let $\hat{\mathbf{G}}_{\mathrm{nys}}$ be the randomized Nyström approximation defined in (11), and let $\hat{\mathbf{G}}_{k}$ be its truncation to rank $k$. Then, for all $k<\ell-1$, the error in expectation satisfyes:
$\mathbb{E}\left[\|{\mathbf{G}-\hat{\mathbf{G}}_{k}}\|_{2}\right]\leq\lambda_{k+1}+\frac{k}{\ell-k-1}\left(\sum_{j>k}\lambda_{j}\right).$
Hence, the Nyström approximation on average achieves the optimal error $\lambda_{k+1}$ for a rank-$k$ approximation, up to an additive term that depends on the tail of the eigenvalues and decreases with $\ell$. This approximation is particularly effective when $\mathbf{G}$ has a strong spectral decay, as it is in our setting. We emphasize that formula (11) is numerically unstable and should not be used directly. A more stable and efficient implementation, which also returns the eigendecomposition of $\hat{\mathbf{G}}_{\mathrm{nys}}$, is described in [60, Algorithm 16] and reported in Algorithm˜1.
Remark 4.2.
The choice of a Gaussian test matrix $\boldsymbol{\Omega}$ over sampling-based methods [2, 6] and other alternatives [5, 60, 63, 37], is motivated by stronger theoretical guarantees [31, 60] and by the fact that in our setting $\mathbf{G}=\mathbf{G}(\boldsymbol{\theta})$ in (8) is accessible only via matrix-vector products, and therefore structured test matrices requiring entrywise access are not advantageous.
1:PSD matrix $\mathbf{G}\in\mathbb{R}^{p\times p}$ (as a function $\mathbf{v}\mapsto\mathbf{G}\mathbf{v}$), rank $\ell$
2:Nyström approximation $\hat{\mathbf{G}}_{\mathrm{nys}}=\mathbf{U}\hat{\boldsymbol{\Lambda}}\mathbf{U}^{\top}$
[…]
Algorithm 1 Randomized Nyström Approximation [31, 60]
[…]
$\mathbf{P}^{-1}=(\hat{\lambda}_{\ell}+\mu)\mathbf{U}(\hat{\boldsymbol{\Lambda}}+\mu\mathbf{I})^{-1}\mathbf{U}^{\top}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\top}).$ (12)
To ensure good performance in expectation, the sampling parameter $\ell$ should be proportional to the effective dimension $d_{\mathrm{eff}}(\mu)=\sum_{i=1}^{p}\lambda_{i}(\mathbf{G})/(\lambda_{i}(\mathbf{G})+\mu)$, which acts as a smoothed count of eigenvalues of $\mathbf{G}$ larger than $\mu$ [31, Theorem 1.1]. When $\mathbf{G}$ exhibits strong spectral decay, $d_{\mathrm{eff}}(\mu)\ll p$, making the randomized Nyström preconditioner particularly effective. The following proposition makes this precise.
Proposition 4.3.
[31, Theorem 5.1] Suppose that the randomized Nyström preconditioner (12) is constructed using Algorithm˜1 with a Gaussian test matrix and $\ell=2\left\lceil 1.5d_{\mathrm{eff}}(\mu)\right\rceil+1$. Then, using $\mathbf{P}^{-1}$ as a preconditioner for $(\mathbf{G}+\mu\mathbf{I})$, results in the condition number bound in expectation:
$\mathbb{E}\left[\kappa_{2}(\mathbf{P}^{-1/2}(\mathbf{G}+\mu\mathbf{I})\mathbf{P}^{-1/2})\right]<28.$
Computational complexity
Each application of the randomized Nyström preconditioner costs $\mathcal{O}(p\ell)$ flops. Beforehand, one must compute the Nyström approximation. Denoting by $\mathrm{C}_{\mathrm{mv}}$ the cost of a matrix-vector product with $\mathbf{G}$, the total cost is $\mathcal{O}(\mathrm{C}_{\mathrm{mv}}\ell+p\ell^{2})$ flops, with memory requirement $\mathcal{O}(p\ell)$. The term $\mathrm{C}_{\mathrm{mv}}\ell$ comes from computing $\mathbf{G}\boldsymbol{\Omega}$, i.e., $\ell$ independent matvecs with $\mathbf{G}$, and is thus embarrassingly parallel. This allows for significant acceleration on multi-core CPUs or GPUs.
4.2 NyströmNGD
We now present our method, NyströmNGD. To compute the regularized natural gradient, we approximately solve $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})\mathbf{d}=\nabla L(\boldsymbol{\theta})$ using the preconditioned conjugate gradient (PCG) method in a matrix-free fashion, with the preconditioner constructed via randomized Nyström. This approach leverages the structure of the problem: matrix-vector products with the Gramian $\mathbf{G}(\boldsymbol{\theta})$ can be computed efficiently via automatic differentiation, while the strong spectral decay commonly observed in practice (see Figure˜1) permits an accurate low-rank approximation, yielding a preconditioner that is both efficient to apply and effective in accelerating convergence. Additionally, the Nyström approximation provides estimates of the top $\ell$ eigenvalues of $\mathbf{G}$, which can be used to adapt the regularization parameter $\mu$ and the rank $\ell$ at each iteration. This section presents the resulting algorithm in full detail.
Adaptation of the regularization parameter $\mu$
The regularization parameter $\mu>0$ is crucial for numerical stability and for mitigating the slow initial convergence typical of quasi-second-order methods like NGD. Prior works proposed heuristics that adapt $\mu$ dynamically [13, 33] or decrease it with the loss [85], so that early iterations mimic gradient descent and gradually shift to NGD near a local minimum. We instead propose a heuristic based on numerical rank of $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$.
Let $\lambda_{1}\geq\dots\geq\lambda_{p}$ denote the eigenvalues of the Gramian $\mathbf{G}(\boldsymbol{\theta})$. We set the regularization parameter as $\mu=\gamma\cdot\epsilon_{\mathrm{mach}}\cdot\lambda_{1}$, with $\gamma=p$ as a reasonable default. This is the threshold commonly employed to compute the numerical rank, i.e., the cut-off under which eigenvalues are considered to be numerically zero [45, 69]. Observe that although $\lambda_{1}$ is not directly available, it can be estimated by $\hat{\lambda}_{1}$ from randomized Nyström. When $\mathbf{G}(\boldsymbol{\theta})$ exhibits strong spectral decay, the expected relative error of this estimate is small even for modest $\ell$ [60, Theorem 14.1]. Finally, we have observed that it can be beneficial to enforce stronger regularization in early iterations, when still far from local minima. This can be achieved with a heuristic of the form $\mu=\max(\gamma\cdot\epsilon_{\mathrm{mach}}\cdot\hat{\lambda}_{1},c\cdot\left\|{\nabla L(\boldsymbol{\theta})}\right\|^{\alpha})$ or $\mu=\max(\gamma\cdot\epsilon_{\mathrm{mach}}\cdot\hat{\lambda}_{1},c\cdot|L(\boldsymbol{\theta})|^{\alpha})$, where the scaling $c$ and the exponent $\alpha$ depend on the problem.
Adaptation of rank parameter $\ell$
The rank parameter $\ell$ controls the computational cost of the preconditioner and should ideally scale with the effective dimension $d_{\mathrm{eff}}(\mu)$ [31, Theorem 5.1]. As $d_{\mathrm{eff}}(\mu)$ is not directly accessible, we estimate $\ell$ adaptively. Following [31, Section 5.4.2], we do it by monitoring the ratio $\hat{\lambda}_{\ell}/\mu$: if it goes above $10$, we double $\ell$ for the next iteration, up to a prescribed maximum $\ell_{\max}$. Otherwise, we reduce $\ell$ to the smallest index satisfying $\hat{\lambda}_{\ell}/\mu>10$, adding a small offset (e.g., $1$ or $2$) to avoid oscillations.
Other algorithmic details and pseudocode
Algorithm˜2 summarizes the proposed algorithm. At each iteration, we construct the randomized Nyström preconditioner (lines 5 to 7) using the matrix-free Gramian. We then apply preconditioned conjugate gradient (PCG) to solve the linear system (6) (line 9). The PCG stopping criterion combines a maximum number of iterations maxit and a relative tolerance $\mathtt{rel\_tol}=\min(\kappa,\left\|{\nabla L(\boldsymbol{\theta}_{k})}\right\|_{2})$, inspired by the Steihaug–Toint truncated CG method [20]. This design avoids overspending computation far from critical points, where NGD steps may underperform compared to GD, and ensures accuracy near local minima. Once the search direction $\mathbf{d}_{k}$ is computed, a line search is performed to update the iterate (lines 10 and 11). In practice, we found that a standard backtracking line-search [83] suffices, without requiring logarithmic grid searches as in [64, 46]. Finally, the rank parameter $\ell$ is updated adaptively (lines 12 to 16).
1:Initial parameter $\boldsymbol{\theta}_{0}\in\mathbb{R}^{p}$, initial rank $\ell_{0}$, maximum rank $\ell_{\max}$, $\gamma\in\mathbb{R}_{+}$ (e.g. $\gamma=p$), CG parameters $\mathtt{maxit},\kappa$ (e.g. $\mathtt{maxit}=20,\kappa=0.1$)
2:Initialize $\ell=\ell_{0}$
3:for $k=0,1,2,\dots$ do
4:  Define $\mathbf{G}_{\mathrm{fun}}=\mathbf{v}\mapsto\mathbf{G}(\boldsymbol{\theta})\mathbf{v}$ via (8) $\triangleright$ Matrix-free Gramian
5:  $\mathbf{U},\hat{\boldsymbol{\Lambda}}]=\mathtt{RandomizedNystrom}(\mathbf{G}_{\mathrm{fun}},\ell) \triangleright$ Randomized Nyström approximation
6:  $\mu=\gamma\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}}$ $\triangleright$ Regularization parameter $\mu$
7:  $\mathbf{P}_{\mathrm{fun}}=\mathbf{v}\mapsto\frac{1}{\hat{\lambda}_{\ell}+\mu}\mathbf{U}(\hat{\boldsymbol{\Lambda}}+\mu\mathbf{I})\mathbf{U}^{\top}\mathbf{v}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\top})\mathbf{v}$ $\triangleright$ Matrix-free preconditioner
8:  $\mathtt{rel\_tol}=\min(\kappa,\left\|{\nabla L(\boldsymbol{\theta}_{k})}\right\|_{2})$ $\triangleright$ Relative tolerance for PCG
9:  $\mathbf{d}_{k}=\texttt{PCG}(\mathbf{G}_{\mathrm{fun}}+\mu\cdot\mathrm{Id},\,\nabla L(\boldsymbol{\theta}_{k}),\,\mathtt{rel\_tol},\,\mathtt{maxit},\,\mathbf{P}_{\mathrm{fun}})$ $\triangleright$ Run PCG
10:  $\alpha_{k}=\texttt{Linesearch}(\boldsymbol{\theta}_{k},-\mathbf{d}_{k})$ $\triangleright$ Linesearch along $\alpha\mapsto\boldsymbol{\theta}_{k}-\alpha\mathbf{d}_{k}$
11:  $\boldsymbol{\theta}_{k+1}=\boldsymbol{\theta}_{k}-\alpha_{k}\mathbf{d}_{k}$
12:  if $\hat{\lambda}_{\ell}>10\mu$ then
13:    $\ell=\min(2\ell,\ell_{\max})$ $\triangleright$ Possibly increase rank parameter $\ell$
14:  else
15:    $\ell=\min\left(\min\{i:\hat{\lambda}_{i}<10\mu\}+1,\,\ell_{\max}\right)$ $\triangleright$ Possibly decrease rank parameter $\ell$
16:  end if
17:end for
Algorithm 2 NyströmNGD
Remark 4.4.
One key advantage of NyströmNGD over standard NGD lies in its reduced and tunable memory requirements. The most memory-intensive step is the Randomized Nyström approximation (line 5, using Algorithm˜1), which at most requires storing two $p\times\ell$ matrices in memory simultaneously. Consequently, the peak memory footprint corresponds to roughly $2\ell$ vectors of length $p$, comparable to the memory complexity of L-BFGS with a history size of $\ell$. This provides explicit control over memory consumption through the parameter $\ell_{\max}$, enabling the method to scale to larger networks. Naturally, smaller values of $\ell_{\max}$ may yield a weaker preconditioner and increase the number of CG iterations, yet the algorithm remains fully operational.
Remark 4.5.
The Randomized Nyström method is not the only viable approach for constructing a symmetric positive definite preconditioner in a matrix-free manner. Alternatives include techniques based on partial Cholesky factorizations. Since the Gramian $\mathbf{G}$ is dense, Incomplete Cholesky is unsuitable; however, one could employ Pivoted Partial Cholesky, either in its deterministic form [43] or its randomized variant [17]. The main limitation of both approaches is that they require access to the full diagonal of $\mathbf{G}$: the deterministic greedy pivoting in [43] and the randomized RPCholesky method in [17] both rely on it. In our context, this involves computing matvecs of $\mathbf{G}$ with all canonical basis vectors. This has the same FLOP cost as constructing the full $\mathbf{G}$, though a lower memory usage. For this reason, we do not pursue these methods further and only include a brief numerical comparison in D.
5 Numerical Experiments
We test NyströmNGD on a range of PDE problems with known analytical solutions. Both strong formulations, discretized with PINNs, and weak formulations, discretized using FEINNs, are considered. Additional numerical experiments are provided in appendix: a sensitivity analysis of NyströmNGD with respect to its key hyperparameters is provided in C, and a brief comparison with alternative preconditioners to randomized Nyström is presented in D.
Comparison of the methods
We compare NyströmNGD against Natural Gradient Descent with an SVD-based pseudoinverse (legend: NGD), matrix-free NGD solved via conjugate gradient without preconditioning (legend: NGD with CG), and the BFGS algorithm [83]. NGD-based methods are run for $500$ or $1000$ iterations, depending on the example, while BFGS is run for $6\times 10^{4}$ or $1.2\times 10^{5}$ iterations. For a fair comparison with NyströmNGD, NGD with CG uses the same relative tolerance as in Algorithm˜2, with a maximum number of iterations set to $\texttt{maxit}+\ell_{\max}$. Given the differing per-iteration costs of the methods, we compare performance in terms of both iteration count and wall-clock time. First-order optimizers (e.g., gradient descent, Adam) and L-BFGS are excluded, as they have been shown to be non-competitive; see Tables˜2 and 1 and [46, 65]. Unless otherwise stated, all experiments use fully connected neural networks with $\tanh$ activation and two hidden layers of width $32$. Each experiment is averaged over 10 runs, with solid lines in the plots indicating medians and shaded regions representing the first and third quantiles.
5.1 PDEs in strong form: PINNs
We begin with PDEs in strong form discretized via Physics-Informed Neural Networks (PINNs) [70, 21].
Computational details
In all PINN experiments, we set in Algorithm˜2 the parameters $\ell_{0}=10$, $\ell_{\max}=500$, $\gamma=p$, $\mathtt{maxit}=20$, and $\kappa=0.1$. For NGD and NGD with CG, we use $\mu=\min(10^{-5},L(\boldsymbol{\theta}))$ as suggested in [85], from which most experiments are adapted. To enforce stronger regularization in early iterations, NyströmNGD employs $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-4}\cdot L(\boldsymbol{\theta})^{2})$. Unless stated otherwise, integrals in the loss and metric are discretized via Monte Carlo quadrature with $N_{\Omega}=3000$ points uniformly sampled in the interior of the domain, and $N_{\Gamma}=800$ points on the boundary. Random integration points are sampled once and fixed throughout training. Errors are always computed using an order of magnitude more quadrature points than during training, sampled independently.
For the PINN formulations, we use JAX [16] and JAXopt [12]. The JAXopt BFGS implementation was optimized for efficiency, as detailed in B, achieving, together with JIT compilation, a $1000\times$ speedup over the timings reported in [64]. The NGD implementation is taken from the public repository of [85]. Integrals in both the loss and the metric are discretized via Monte Carlo quadrature. All experiments are conducted in double precision on a single NVIDIA RTX 4090 GPU with 24 GB of memory.
A comparison with other optimizers
As already noted, first-order optimizers and L-BFGS have been shown to be non-competitive compared to NGD and BFGS [46, 65]. To experimentally confirm this, we include a brief comparison of NyströmNGD and BFGS with other state-of-the-art algorithms: L-BFGS [83], Adam [51], Shampoo [35, 4], and Muon [47]. For L-BFGS, Adam, and Muon, we use the Optax implementation [26]; for Shampoo, we use the implementation by Google Research. Results are reported in Tables˜1 and 2 for the three problems discussed later in detail.
Our proposed method consistently outperforms the other optimizers in terms of final accuracy, even when they are given a larger computational time budget. Moreover, as shown in Table˜2, NyströmNGD reaches the same accuracy as other methods in substantially less time. These results clearly indicate that the best runner-up to NyströmNGD is the BFGS method, which is therefore the only one included in the subsequent comparisons.
Table 1: PDEs in strong form discretized via PINNs. Final relative $\mathrm{H}^{1}$ error, for different optimizers.
Poisson Heat Kovasznay
NyströmNGD $2.92\text{\times}{10}^{-05}$ $2.34\text{\times}{10}^{-05}$ $3.01\text{\times}{10}^{-06}$
BFGS $1.88\text{\times}{10}^{-04}$ $34\text{\times}{10}^{-04} 1.1$$10\text{\times}{10}^{-05}$
L-BFGS $8.28\text{\times}{10}^{-04}$ $4.42\text{\times}{10}^{-04}$ $7.93\text{\times}{10}^{-05}$
Adam $2.19\text{\times}{10}^{-02}$ $6.77\text{\times}{10}^{-03}$ $4.18\text{\times}{10}^{-03}$
Shampoo $8.44\text{\times}{10}^{-02}$ $2.52\text{\times}{10}^{-02}$ $8.84\text{\times}{10}^{-01}$
Muon $2.84\text{\times}{10}^{-02}$ $7.78\text{\times}{10}^{-03}$ $1.48\text{\times}{10}^{-03}$
Table 2: PDEs in strong form discretized via PINNs. Total training time (in seconds) for different optimizers; NyströmNGD’s time to reach the same accuracy is shown in parentheses.
Poisson Heat Kovasznay
NyströmNGD 86.1 335.4 485.2
BFGS 81.7 (11.1) 304.6 (22.0) 363.6 (226.5)
L-BFGS 295.4 (9.7) 685.7 (19.4) 476.3 (189.0)
Adam 228.5 (5.3) 911.4 (13.7) 504.7 (97.9)
Shampoo 503.5 (1.9) 1052.0 (6.9) 766.6 (0.8)
Muon 277.5 (4.6) 951.9 (12.7) 635.2 (127.3)
5.1.1 Poisson problem in 3D
We consider the Poisson problem on the 3D unit cube
$\begin{cases}-\Delta u=f&\Omega=(0,1)^{3},\\ u=g_{D}&\partial\Omega,\end{cases}$ (13)
with exact solution $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$, as in [85]. Problem (13) is linear and is considered in strong form. Using the notation of Section˜2.2.1, the function spaces are $\mathrm{V}=\mathrm{H}^{2}(\Omega)$, $\mathrm{X}=\mathrm{L}^{2}(\Omega)$, and $\mathrm{Y}=\mathrm{L}^{2}(\partial\Omega)$, with operators
$\mathcal{F}(u)=-\Delta u,\qquad\mathcal{B}(u)=\left.{u}\right|_{\partial\Omega}.$
The corresponding least-squares loss is
$E(u)=\frac{1}{2}\left\|{\Delta u+f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(\partial\Omega)}^{2},$ (14)
and the induced metric reads
${g}\left({u},\,{v}\right)=\left({\Delta u},\,{\Delta v}\right)_{\mathrm{L}^{2}(\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\partial\Omega)}=\frac{1}{2}\int_{\Omega}\Delta u\,\Delta v\,\mathrm{d}\mathbf{x}+\frac{1}{2}\int_{\partial\Omega}u\,v\,\mathrm{d}\mathbf{s}.$ (15)
Remark 5.1.
The metric (15) is coercive only in the $\mathrm{H}^{1/2}(\Omega)$ norm, but not in $\mathrm{H}^{2}(\Omega)$. This is due to the use of the $\mathrm{L}^{2}(\partial\Omega)$ norm for the Dirichlet boundary term. Coercivity in $\mathrm{H}^{2}(\Omega)$ can be restored by enforcing the boundary conditions strongly [85, Remark 6] or by penalizing in the $\mathrm{H}^{1/2}(\partial\Omega)$ norm [14]. However, in line with standard PINN practice, we retain the $\mathrm{L}^{2}(\partial\Omega)$ penalty for simplicity.
Results
As shown in Figure˜2, NyströmNGD yields the best results in both final accuracy and computational efficiency, converging in around $100$ iterations. The Nyström rank parameter $\ell$ increases rapidly during early iterations, consistent with Figure˜3, and then stabilizes well below the maximum allowed rank $\ell_{\max}=500$. This confirms the effectiveness of the adaptive rank selection.
NGD initially converges faster per iteration but quickly plateaus. According to [85], it would require roughly $5000$ iterations to reach the accuracy achieved by NyströmNGD, resulting in a substantially higher total cost. Importantly, NyströmNGD not only lowers computational cost but also improves final accuracy compared to standard NGD. While we do not have a definitive explaination, this behavior may stem from enhanced numerical stability or from the stochastic nature of the Nyström approximation, which can help escape poor local minima, analogous to how SGD can outperform full-batch GD in classical machine learning [53, 48, 40]. NGD with CG performs similarly to NGD in iteration count but is considerably slower due to the large number of CG steps needed to approximate the Gramian inverse, underscoring the crucial role of preconditioning. BFGS is not competitive in iteration count; although its low per-iteration cost yields faster initial progress in wall-clock time, it ultimately stagnates at a higher error.
(a) Relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $500$ iterations)
(b) Relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time.
(c) Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 2: Convergence for the 3D Poisson equation discretized via PINNs.
Figure 3: 3D Poisson equation discretized via PINNs and trained with NGD using an SVD-based pseudoinverse. Left: decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. Right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
5.1.2 Scalability
Although an MLP with two hidden layers of 32 neurons each was already sufficient to achieve a relative accuracy on the order of $10^{-5}$, limited primarily by quadrature error, we also assess the scalability of NyströmNGD. As discussed in Remark˜4.4, NyströmNGD provides explicit memory control via the parameter $\ell_{\max}$, enabling its application to significantly larger networks than standard NGD.
We perform both width ($32\to 256$) and depth ($2\to 6$) sweeps for the 3D Poisson problem. We set $\ell_{\max}=500$, except for the networks with $128$ and $256$ neurons per layer, where we use $\ell_{\max}=300$ and $\ell_{\max}=150$, respectively. We compare NyströmNGD against BFGS and L-BFGS. Results are summarized in Tables˜3, 4, 5 and 6.
NyströmNGD scales effectively to networks substantially larger than those trainable with standard NGD, consistently outperforming both BFGS and L-BFGS in terms of final accuracy ( Tables˜3 and 5) and total runtime ( Tables˜4 and 6). As expected, decreasing $\ell_{\max}$ leads to a weaker preconditioner and an increased number of CG iterations, as shown in Tables˜7 and 8, yet NyströmNGD continues to provide the best trade-off between computational cost and accuracy. For instance, on the largest tested network (width 256, 67k params), BFGS runs out of memory on our GPU. In contrast, NyströmNGD, with memory usage comparable to L-BFGS, reaches a lower error with a $7\times$ speedup for the same accuracy. These findings underscore the improved scalability of NyströmNGD over standard NGD.
Table 3: Width sweep ($32\to 256$) on the 3D Poisson problem discretized via PINNs. Final relative $\mathrm{H}^{1}$ errors.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{64\times 2}$ (4.5k) $\boldsymbol{128\times 2}$ (17k) $\boldsymbol{256\times 2}$ (67k)
NyströmNGD $2.37\text{\times}{10}^{-05}$ $1.28\text{\times}{10}^{-05}$ $1.83\text{\times}{10}^{-05}$ $7.53\text{\times}{10}^{-05}$
BFGS $1.86\text{\times}{10}^{-04}$ $7.07\text{\times}{10}^{-05}$ $5.59\text{\times}{10}^{-05}$ Out Of Memory
L-BFGS $1.32\text{\times}{10}^{-03}$ $7.12\text{\times}{10}^{-04}$ $6.00\text{\times}{10}^{-04}$ $4.57\text{\times}{10}^{-04}$
Table 4: Width sweep ($32\to 256$) on the 3D Poisson problem discretized via PINNs. Total optimization times (in seconds); NyströmNGD’s time to reach the same accuracy is shown in parentheses.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{64\times 2}$ (4.5k) $\boldsymbol{128\times 2}$ (17k) $\boldsymbol{256\times 2}$ (67k)
NyströmNGD 38.2 94.6 247.1 541.6
BFGS 71.4 (15.4) 125.3 (20.6) 678.5 (51.2) Out Of Memory
L-BFGS 204.8 (11.2) 260.9 (15.7) 378.0 (34.4) 762.5 (111.8)
Table 5: Depth sweep ($2\to 6$) on the 3D Poisson problem discretized via PINNs. Final relative $\mathrm{H}^{1}$ errors.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{32\times 3}$ (2.3k) $\boldsymbol{32\times 4}$ (3.3k) $\boldsymbol{32\times 5}$ (4.4k) $\boldsymbol{32\times 6}$ (5.4k)
NyströmNGD $37\text{\times}{10}^{-05} 2.2$$21\text{\times}{10}^{-05}$ $2.59\text{\times}{10}^{-05}$ $4.35\text{\times}{10}^{-05}$ $2.10\text{\times}{10}^{-05}$
BFGS $1.86\text{\times}{10}^{-04}$ $1.42\text{\times}{10}^{-04}$ $1.25\text{\times}{10}^{-04}$ $9.52\text{\times}{10}^{-05}$ $8.70\text{\times}{10}^{-05}$
L-BFGS $32\text{\times}{10}^{-03} 1.1$$17\text{\times}{10}^{-03}$ $1.21\text{\times}{10}^{-03}$ $50\text{\times}{10}^{-03} 1.1$$18\text{\times}{10}^{-03}$
Table 6: Depth sweep ($2\to 6$) on the 3D Poisson problem discretized via PINNs. Total optimization times (in seconds); NyströmNGD’s time to reach the same accuracy is shown in parentheses.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{32\times 3}$ (2.3k) $\boldsymbol{32\times 4}$ (3.3k) $\boldsymbol{32\times 5}$ (4.4k) $\boldsymbol{32\times 6}$ (5.4k)
NyströmNGD 38.2 59.4 80.6 87.5 121.7
BFGS 71.4 (15.4) 107.5 (13.3) 153.5 (22.8) 199.8 (29.5) 247.4 (24.6)
L-BFGS 204.8 (11.2) 248.5 (10.6) 291.2 (17.0) 341.1 (23.5) 382.1 (17.2)
Table 7: Width sweep ($32\to 256$) on the 3D Poisson problem discretized via PINNs. Average Nyström rank parameter $\ell$ (with $\ell_{\max}$ in parentheses) and number of CG iterations for NyströmNGD.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{64\times 2}$ (4.5k) $\boldsymbol{128\times 2}$ (17k) $\boldsymbol{256\times 2}$ (67k)
[…]
CG it. 13 17 22 38
Table 8: Depth sweep ($2\to 6$) on the 3D Poisson problem discretized via PINNs. Average Nyström rank parameter $\ell$ (with $\ell_{\max}$ in parentheses) and number of CG iterations for NyströmNGD.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{32\times 3}$ (2.3k) $\boldsymbol{32\times 4}$ (3.3k) $\boldsymbol{32\times 5}$ (4.4k) $\boldsymbol{32\times 6}$ (5.4k)
[…]
CG it. 13 15 16 16 17
5.1.3 Heat equation in 3+1D
We consider the heat equation
$\begin{cases}\partial_{t}u-\Delta u=f&\text{in }I\times\Omega=(0,1)\times(0,1)^{3},\\ u=g_{D}&\text{on }I\times\partial\Omega,\\ u(0,\cdot)=u_{0}&\text{in }\Omega,\end{cases}$ (16)
with exact solution $u(t,x,y,z)=\big(\cos(\pi x)+\cos(\pi y)+\cos(\pi z)\big)e^{-\pi^{2}t/4}$, as in [85]. We consider Equation˜16 in strong form, with function spaces $\mathrm{V}=\mathrm{L}^{2}(I,\mathrm{H}^{2}(\Omega))\cap\mathrm{H}^{1}(I,\mathrm{L}^{2}(\Omega))$, $\mathrm{X}=\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))$, $\mathrm{Y}=\mathrm{L}^{2}(I,\mathrm{L}^{2}(\partial\Omega))\times\mathrm{L}^{2}(\Omega)$, and operators
$\mathcal{F}(u)=\partial_{t}u-\Delta u,\qquad\mathcal{B}(u)=\big(\left.{u}\right|_{I\times\partial\Omega},\,u(0,\cdot)\big).$
The corresponding least-squares loss reads
$E(u)=\frac{1}{2}\left\|{\partial_{t}u-\Delta u-f}\right\|_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\partial\Omega))}^{2}+\frac{1}{2}\left\|{u(0,\cdot)-u_{0}}\right\|_{\mathrm{L}^{2}(\Omega)}^{2},$ (17)
with induced metric
${g}\left({u},\,{v}\right)=\left({\partial_{t}u-\Delta u},\,{\partial_{t}v-\Delta v}\right)_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}+\left({u(0,\cdot)},\,{v(0,\cdot)}\right)_{\mathrm{L}^{2}(\Omega)}.$ (18)
Considerations on the coercivity of the metric analogous to Remark˜5.1 apply here; see [85, Section 4.5] for details.
Results
As shown in Figure˜4, NyströmNGD again provides the best accuracy at a reduced computational cost. All other methods saturate at a lower accuracy, despite being given a higher computational budget. Comparing with [85], NGD would not reach the final accuracy of NyströmNGD even when given a larger computational budget of $5000$ iterations. The Nyström rank parameter $\ell$ increases rapidly during early iterations and then stabilizes below $\ell_{\max}=500$, but at a higher value than in the Poisson example, consistent with the slightly slower singular value decay observed in Figure˜5.
(a) Relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations).
(b) Relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time.
(c) Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 4: Convergence for the Heat equation in 3+1D discretized via PINNs.
Figure 5: Heat equation in 3+1D discretized via PINNs and trained with NGD using the SVD-based pseudoinverse. Left: decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. Right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
5.1.4 Kovasznay Flow (2D steady state Navier–Stokes flow)
We consider the two-dimensional steady Navier–Stokes flow as originally described in [54], with Reynolds number $\mathrm{Re}=40$, following the setup in [46, Section 4.1]. The system of PDEs reads
$\begin{cases}-\nu\Delta u+(u\cdot\nabla)u+\nabla p=f&\Omega=(-\frac{1}{2},1)\times(-\frac{1}{2},\frac{3}{2}),\\ \nabla\cdot u=0&\Omega,\\ u=g_{D}&\partial\Omega,\end{cases}$ (19)
with exact solution
$\begin{split}u(x,y)&=\begin{bmatrix}1-e^{\lambda x}\cos(2\pi y)\\ \frac{\lambda}{2\pi}e^{\lambda x}\sin(2\pi y)\end{bmatrix},\\ p(x,y)&=\frac{1}{2}(1-e^{2\lambda x})-c,\end{split}$
where $\lambda=\frac{1}{2\nu}-\sqrt{\frac{1}{4\nu^{2}}+4\pi^{2}}$, $\nu=\frac{1}{\mathrm{Re}}=\frac{1}{40}$, and $c$ is chosen such that $\int_{\Omega}p=0$. Following [46], we consider the function spaces $\mathrm{V}=\mathrm{H}^{2}(\Omega)\times\{p\in\mathrm{H}^{1}(\Omega):\int_{\Omega}p=0\}$ and $\mathrm{X}=\mathrm{L}^{2}(\Omega)^{2}\times\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\partial\Omega)$. Problem (19) is written in strong form as a nonlinear least-squares problem with residual
$\mathcal{R}(u,p)=\begin{bmatrix}-\nu\Delta u+(u\cdot\nabla)u+\nabla p-f\\ \nabla\cdot u\\ u-g_{D}\end{bmatrix}$, (20)
and corresponding energy functional
$E(u,p)=\frac{1}{2}\left\|{-\nu\Delta u+(u\cdot\nabla)u+\nabla p-f}\right\|^{2}_{\mathrm{L}^{2}(\Omega)^{2}}+\frac{1}{2}\left\|{\nabla\cdot u}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{1}{2}\left\|{u-g_{D}}\right\|^{2}_{\mathrm{L}^{2}(\partial\Omega)}.$ (21)
As a metric, we consider the Gauss–Newton metric
${g_{(u,p)}}\left({(v,h)},\,{(w,t)}\right)=\left({\mathrm{D}\mathcal{R}(u,p)[(v,h)]},\,{\mathrm{D}\mathcal{R}(u,p)[(w,t)]}\right).$
Since the residual (20) is nonlinear, the metric is point-dependent. We refer to [46] for a discussion on coercivity. To discretize (19), we represent $(u,p)=(u_{\boldsymbol{\theta}_{1}},p_{\boldsymbol{\theta}_{2}})$ as two neural networks, with parameters combined into a single vector $\boldsymbol{\theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}]$.
Results
In this example, NyströmNGD again achieves the best accuracy in fewer than $1000$ iterations, as shown in Figure˜6. However, compared to the previous examples, the Nyström rank parameter $\ell$ exhibits a pronounced peak during intermediate iterations, reaching the maximum rank $\ell_{\max}=500$ before stabilizing around $400$. This behavior indicates that the Gramian is less suited to low-rank approximation in this problem, which explains the smaller differences in final accuracy and computational time between NyströmNGD and BFGS. Nevertheless, NyströmNGD still requires substantially fewer iterations to converge.
(a) Relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations (first $1000$ iterations.)
(b) Relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. wall-clock time.
(c) Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 6: Convergence for the Kovasnay Flow discretized via PINNs.
Figure 7: Kovasnay Flow discretized via PINNs and trained with NGD using SVD-based pseudoinverse. Left: decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. Right: the relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations.
5.2 PDEs in weak form: FEINNs
We now consider variational formulations of PDEs using FEINNs, as defined in Section˜2.3.1, to approximate the solution.
Computational details
For all experiments with variational formulations, the parameters in Algorithm˜2 are set as $\ell_{0}=10$, $\ell_{\max}=500$, $\gamma=30$, $\mathtt{maxit}=20$, and $\kappa=0.1$. Neural networks are interpolated onto a finite element space of degree $d=4$ over a uniform quadrilateral mesh with mesh size $h=1/30$, yielding approximately $14,500$ degrees of freedom. Integrals in the loss and metric are evaluated by interpolation onto the FE space and Gauss-Jacobi quadrature of order $2d=8$. Errors are computed with a Gauss-Jacobi quadrature of order $4d=16$. We use the Julia packages Lux.jl [67] for neural networks and Optim.jl [62] for optimization. All experiments are conducted in double precision on a single NVIDIA RTX 4090 GPU with 24 GB of memory.
5.2.1 Poisson problem in 2D – Energy minimization (or Deep Ritz like) formulation
We consider the Poisson problem on the 2D unit square
$\begin{cases}-\Delta u=f&\Omega=(0,1)^{2},\\ u=g_{D}&\partial\Omega,\end{cases}$ (22)
where $f\in\mathrm{H}^{-1}(\Omega)$ and $g_{D}\in\mathrm{H}^{\frac{1}{2}}(\partial\Omega)$. As exact solution we take $u(x,y)=\sin(\pi x)\sin(\pi y)$.
All equalities in (22) are understood in the variational sense. That is, given a lifting $\tilde{g}_{D}\in\mathrm{H}^{1}(\Omega)$ such that $\left.{\tilde{g}_{D}}\right|_{\partial\Omega}=g_{D}$, we seek $u_{0}\in\mathrm{H}^{1}_{0}(\Omega)$ satisfying
$a(u_{0},v)=\int_{\Omega}fv\,\mathrm{d}\mathbf{x}-a(\tilde{g}_{D},v)\quad\forall v\in\mathrm{H}^{1}_{0}(\Omega),\qquad\text{where}\quad a(u,v)=\int_{\Omega}\nabla u\cdot\nabla v\,\mathrm{d}\mathbf{x}.$ (23)
The solution is then $u=u_{0}+\tilde{g}_{D}$. The discretization spaces are $\mathrm{V}=\mathrm{H}^{1}_{0}(\Omega)$, $\mathrm{X}=\mathrm{V}^{\prime}=\mathrm{H}^{-1}(\Omega)$, with linear operator
$\mathcal{F}:\mathrm{H}^{1}_{0}(\Omega)\to\mathrm{H}^{-1}(\Omega),\qquad\left\langle{\mathcal{F}u},\,{v}\right\rangle_{\mathrm{H}^{-1}\times\mathrm{H}^{1}_{0}}=a(u,v).$
Since $a$ is symmetric and coercive on $\mathrm{V}\times\mathrm{V}$, we adopt an energy minimization formulation, akin to the Deep Ritz method:
$\min_{u_{0}\in\mathrm{H}^{1}_{0}(\Omega)}E(u_{0})\coloneqq\int_{\Omega}\frac{1}{2}\left\|{\nabla u_{0}}\right\|^{2}-fu_{0}+\nabla\tilde{g}_{D}\cdot\nabla u_{0}\,\mathrm{d}\mathbf{x},$ (24)
with corresponding metric $g=a$.
Computational details
We run NGD-based methods for $500$ iterations and BFGS for $1.2\times 10^{5}$ iterations. Since the loss can become negative, we set $\mu=10^{-9}$ for NGD and NGD with CG, and $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-9})$ for NyströmNGD.
Results
NyströmNGD and BFGS achieve comparable final accuracy, as shown in Figure˜8, but NyströmNGD requires a fraction of the iterations and is approximately $10\times$ faster in wall-clock time. This efficiency stems from the strong low-rank approximability of the Gramian, evident in Figure˜9, which leads the Nyström rank parameter $\ell$ to stabilize around $80$, far below the parametric dimension $p$. NGD and NGD with CG plateau at lower accuracy and are not competitive in iterations or runtime.
(a) Relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations).
(b) Relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time.
(c) Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 8: Convergence for the Poisson equation in 2D discretized via FEINNs and energy minimization.
Figure 9: Poisson equation in 2D discretized via FEINNs and energy minimization, trained with NGD using an SVD-based pseudoinverse. Left: decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. Rigth: the relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
5.2.2 Diffusion-reaction-transport problem in 2D
We consider the diffusion-reaction-transport problem
$\begin{dcases}-\nabla\cdot(\kappa\nabla u)+\boldsymbol{\beta}\cdot\nabla u+\sigma u=F&\Omega=(0,1)^{2},\\ u=g&\Gamma_{D}\subset\partial\Omega,\\ \kappa\nabla u\cdot\mathbf{n}=\psi&\Gamma_{N}\subset\partial\Omega,\end{dcases}$ (25)
where $\kappa(x,y)=2+\sin(x+2y)$, $\boldsymbol{\beta}(x,y)=\begin{bmatrix}\sqrt{x-y^{2}+5}&\sqrt{y-x^{2}+5}\end{bmatrix}^{\top}$, and $\sigma(x,y)=e^{\frac{x}{2}-\frac{y}{3}}+2$. The source term $f$ is chosen so that the exact solution is
$u(x,y)=\sin(3.2x(x-y))\cos(4.3y+x)+\sin(4.6(x+2y))\cos(2.6(y-2x)),$
as in [10, 8]. The Dirichlet boundary $\Gamma_{D}$ consists of the left and right sides of the domain, while the Neumann boundary $\Gamma_{N}$ corresponds to the top and bottom edges.
The function spaces are $\mathrm{V}=\mathrm{H}^{1}_{0,\Gamma_{D}}(\Omega)$, $\mathrm{X}=\mathrm{V}^{\prime}=\mathrm{H}^{-1}_{\Gamma_{D}}(\Omega)$, and the variational formulation of (25) is defined as $\mathcal{F}u=f$ where
$\begin{split}\left\langle{\mathcal{F}u},\,{v}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}}&=a(u,v)=\int_{\Omega}(\kappa\nabla u)\cdot\nabla v+(\boldsymbol{\beta}\cdot\nabla u)\,v+\sigma\,u\,v\,\,\mathrm{d}\mathbf{x}\\ \left\langle{f},\,{v}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}}&=\int_{\Omega}F\,v\,\,\mathrm{d}\mathbf{x}+\int_{\Gamma_{N}}\psi\,v\,\,\mathrm{d}\mathbf{s}\end{split}$
Since $\mathcal{F}$ is in general not symmetric positive definite, we adopt the least-squares formulation
$\min_{u\in\mathrm{H}^{1}_{0}(\Omega)}E(u)\coloneqq\frac{1}{2}\left\|{\mathcal{F}u-f}\right\|_{\mathrm{H}^{-1}_{\Gamma_{D}}}^{2}=\left\langle{\mathcal{F}u-f},\,{\mathcal{R}_{\mathrm{V}}^{-1}(\mathcal{F}u-f)}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}},$ (26)
with associated metric ${g}\left({v},\,{w}\right)\coloneqq\left\langle{\mathcal{F}w},\,{\mathcal{R}_{\mathrm{V}}^{-1}\mathcal{F}v}\right\rangle$. Here, $\mathcal{R}_{\mathrm{V}}^{-1}=(-\Delta)^{-1}:\mathrm{H}^{-1}_{\Gamma_{D}}\to\mathrm{H}^{1}_{0,\Gamma_{D}}$ denotes the inverse Riesz operator, which involves solving a Poisson equation.
When discretizing (26) and the corresponding metric with FEINNs, applying the discrete Riesz operator $\mathcal{R}_{\mathrm{V}_{h}}^{-1}$ corresponds to solving a Poisson problem with FEM. To reduce computational cost, one could instead use a spectrally equivalent approximation, such as a few cycles of a geometric multigrid preconditioner, as done in [8]. However, for simplicity, we solve the Poisson problem directly using FEM, computing a Cholesky factorization of the Laplacian matrix once.
Computational details
We run NGD-based methods for $1000$ iterations and BFGS for $1.2\times 10^{5}$ iterations. For NGD and NGD with CG we use $\mu=\min(10^{-5},L(\boldsymbol{\theta}))$ as suggested in [85]. For NyströmNGDwe set $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-8}\cdot L(\boldsymbol{\theta})^{2})$.
Results
As illustrated in Figure˜10, NyströmNGD reaches convergence in few iterations and achieves higher accuracy compared to other NGD methods. The gap in convergence speed relative to BFGS is substantial in both iterations and wall-clock time. The smaller difference in wall-clock time between NyströmNGD and BFGS, compared to previous variational examples, is due to the higher Nyström rank parameter $\ell$ at which NyströmNGD stabilizes. Although $\ell$ remains much lower than the parametric dimension $p$, it is higher than in earlier examples, reflecting the slower, yet still significant, decay of the singular values of the Gramian shown in Figure˜11.
(a) Relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations).
(b) Relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time.
(c) Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 10: Convergence for the diffusion-reaction-transport equation (25) in 2D discretized via FEINNs.
Figure 11: Diffusion-reaction-transport equation in 2D discretized via FEINNs, trained with NGD using an SVD-based pseudoinverse. Left: decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. Right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
6 Discussion and Conclusion
We introduced NyströmNGD, a novel optimization algorithm that uses randomized Nyström preconditioning to accelerate the inner conjugate gradient (CG) solver in matrix-free natural gradient descent (NGD). This approach significantly reduces the cost of the inner solver, resulting in faster convergence in both iterations and wall-clock time. Moreover, our work extends matrix-free NGD to a broader class of metrics relevant to PDE-related problems. We validate NyströmNGD on various PDEs problems, both in strong form discretized with PINNs and in variational form discretized with FEINNs, demonstrating that it consistently outperforms previous NGD methods by achieving higher accuracy at lower computational cost.
The use of randomized Nyström for preconditioning builds on the empirically observed low-rank structure of the Gramian. Our results confirm that in many cases the Gramian admits strong low-rank compression, though this is not always guaranteed and warrants further analysis. Nevertheless, the core idea of preconditioning the inner CG solver remains applicable, and we plan to explore alternative Gramian approximations for use as preconditioners. Notably, NyströmNGD not only reduces the computational cost but also improves final accuracy over standard NGD. This may be due to improved numerical stability or the stochastic nature of the Nyström approximation. This aspect merits further investigation. Finally, NyströmNGD is especially promising in the context of neural network-based solvers for variational formulations of PDEs, where the choice of metric can be guided by the underlying variational structure. Our results in Section˜5.2 support the potential of the proposed approach and motivate future research in this direction.
CRediT authorship contribution statement
Ivan Bioli: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Writing – Original Draft, Writing – Review & Editing, Visualization
Carlo Marcati: Conceptualization, Methodology, Validation, Formal analysis, Writing – Original Draft, Writing – Review & Editing, Supervision
Giancarlo Sangalli: Conceptualization, Methodology, Validation, Formal analysis, Resources, Writing – Original Draft, Writing – Review & Editing, Supervision, Project administration, Funding acquisition
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
No data were used for the research described in the article. The code used to reproduce the findings of this paper is available from the corresponding author upon request.
Acknowledgments
This work has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie Action MSCA-DN-101119556 (IN-DEEP). The authors are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). C. Marcati and G. Sangalli acknowledge support from the Italian MUR through the PRIN 2022 PNRR project NOTES (No. P2022NC97R). G. Sangalli further acknowledges support from the Italian MUR through PNRR-M4C2-I1.4-NC-HPC-Spoke6. C. Marcati further acknowledges the PRIN 2022 project ASTICE (No. 202292JW3F).
Appendix A Choice of the Metric
The link with operator preconditioning in Section˜3.2.1 makes it evident that the choice of metric has a crucial influence on the convergence speed of NGD. In some cases, the problem’s structure naturally suggests a metric, as in Variational MonteCarlo [68, 42, 65] or Sobolev gradient flows [41, 19]. In general, the link between NGD and Riemannian gradient descent [64] suggests selecting the metric $g$ to reduce the condition number of the Riemannian Hessian [15]. Indeed, it can be shown that NGD corresponds to Riemannian gradient descent [15] on the manifold $(\mathrm{V},g)$, up to the $g$-orthogonal projection of the Riemannian gradient onto the model’s generalized tangent space $\mathrm{T}_{\boldsymbol{\theta}}\mathcal{N}$, and a second-order correction; see [65, Theorem 1]. More precisely, the NGD update $\boldsymbol{\theta}_{k+1}=\boldsymbol{\theta}_{k}-\alpha_{k}\mathbf{G}(\boldsymbol{\theta}_{k})^{\dagger}\nabla L(\boldsymbol{\theta}_{k})$ satisfies
$u_{\boldsymbol{\theta}_{k+1}}=u_{\boldsymbol{\theta}_{k}}-\alpha_{k}\Pi_{\boldsymbol{\theta}_{k}}^{g}\left(\mathrm{grad}^{g}E(u_{\boldsymbol{\theta}_{k}})\right)+\varepsilon_{k},$
where $\Pi_{\boldsymbol{\theta}_{k}}^{g}$ denotes the $g$-orthogonal projection onto the generalized tangent space $\mathrm{T}_{\boldsymbol{\theta}_{k}}\mathcal{N}$, and $\varepsilon_{k}=\mathcal{O}\big(\alpha_{k}^{2}\left\|{\mathbf{G}(\boldsymbol{\theta}_{k})^{\dagger}\nabla L(\boldsymbol{\theta}_{k})}\right\|^{2}\big)$. A principled strategy for selecting the metric $g$ is thus to adopt criteria from Riemannian optimization, choosing $g$ to reduce the condition number of the Riemannian Hessian [15, 1]. Although convergence may differ from standard Riemannian gradient descent due to the projection onto $\mathrm{T}_{\boldsymbol{\theta}}\mathcal{N}$, whose effect depends on the geometry of $\mathcal{N}$ and the parametrization, this strategy can be expected to remain effective if the ansatz $\mathcal{N}$ is sufficiently expressive.
A.1 Linear problems
Setting A.1.
Let $\mathrm{V},\mathrm{W}$ be Hilbert spaces, and let $\mathcal{A}:\mathrm{V}\to\mathrm{W}^{\prime}$ be a linear isomorphism. Given $f\in\mathrm{W}^{\prime}$ we aim at solving
$\mathcal{A}u=f\in\mathrm{W}^{\prime}.$ (27)
A natural approach is to consider the least-squares problem
$\inf_{u\in\mathrm{V}}E(u)\coloneqq\frac{1}{2}\left\|{\mathcal{A}u-f}\right\|_{\mathrm{W}^{\prime}}^{2}.$ (28)
Choosing on $\mathrm{V}=\mathrm{V}$ the (point-independent) metric
${g}\left({v},\,{w}\right)\coloneqq\left({\mathcal{A}v},\,{\mathcal{A}w}\right)_{\mathrm{W}^{\prime}},$ (29)
yields $\mathrm{Hess}^{g}E=\mathrm{Id}$. Riemannian gradient descent on $(\mathrm{V},g)$ thus converges in a single iteration, as the Riemannian gradient is given by $\mathrm{grad}^{g}E(u)=u-\mathcal{A}^{-1}f$, i.e., computing it is equivalent to solving (27). NGD does not require applying $\mathcal{A}^{-1}$: computing the Gramian $\left[\mathbf{G}(\boldsymbol{\theta})\right]_{ij}=\left({\mathcal{A}\partial_{\theta_{i}}u_{\boldsymbol{\theta}}},\,{\mathcal{A}\partial_{\theta_{j}}u_{\boldsymbol{\theta}}}\right)_{\mathrm{W}^{\prime}}$ involves only applications of $\mathcal{A}$.
Remark A.2.
We formulate the linear problem from $\mathrm{V}$ to the dual of $\mathrm{W}$ because one of our target applications involves variational formulations of PDEs, which naturally take the form (27). For PINNs, which are based on strong formulations, $\mathrm{W}$ is typically a (product of) $\mathrm{L}^{2}$ spaces, so that $\mathrm{W}^{\prime}\cong\mathrm{W}$ and the formulation recovers that of [85], which does not employ the dual.
Remark A.3.
In Setting A.1, it is typically straightforward to approximate numerically the inner product $\left({\cdot},\,{\cdot}\right)_{\mathrm{W}}$ and the duality pairing $\left\langle{\cdot},\,{\cdot}\right\rangle_{\mathrm{W}^{\prime}\times\mathrm{W}}$, but not the dual inner product $\left({\cdot},\,{\cdot}\right)_{\mathrm{W}^{\prime}}$. Let $\mathcal{R}_{\mathrm{W}}:\mathrm{W}\to\mathrm{W}^{\prime}$ denote the Riesz map, defined by
$\left\langle{\mathcal{R}_{\mathrm{W}}v},\,{w}\right\rangle_{\mathrm{W}^{\prime}\times\mathrm{W}}=\left({v},\,{w}\right)_{\mathrm{W}}\qquad\forall v,w\in\mathrm{W}.$
Then, for $\ell,h\in\mathrm{W}^{\prime}$, we have
$\left({\ell},\,{h}\right)_{\mathrm{W}^{\prime}}=\left\langle{\ell},\,{\mathcal{R}_{\mathrm{W}}^{-1}h}\right\rangle_{\mathrm{W}^{\prime}\times\mathrm{W}}.$
For example, when $\mathrm{W}=\mathrm{H}^{1}_{0}(\Omega)$, the dual space is $\mathrm{W}^{\prime}=\mathrm{H}^{-1}(\Omega)$, and $\mathcal{R}_{\mathrm{W}}=\Delta$, the Laplace operator. Computing the dual inner product thus requires applying $(-\Delta^{-1})$, i.e., solving a (discretized) Laplace problem. In contrast, for $\mathrm{W}=\mathrm{L}^{2}(\Omega)$, one has $\mathrm{W}^{\prime}\cong\mathrm{W}$ and $\mathcal{R}_{\mathrm{W}}=\mathrm{Id}$.
Symmetric positive definite linear operators
If $\mathrm{V}=\mathrm{W}$ and $\mathcal{A}$ is symmetric positive definite, one can endow $\mathrm{V}$ with the $\mathcal{A}$-induced inner product
$\left({v},\,{w}\right)_{\mathcal{A}}=\left\langle{\mathcal{A}v},\,{w}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}\qquad\forall v,w\in\mathrm{V}.$
This corresponds to choosing $\mathcal{R}_{\mathrm{V}}=\mathcal{A}$, and the loss (28) becomes
$E(u)=\frac{1}{2}\left\langle{\mathcal{A}u},\,{u}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}-\left\langle{f},\,{u}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}+\mathrm{const},$ (30)
where the constant is irrelevant for gradient-based optimization. This is precisely the variational form used in the Deep Ritz method [30]. The corresponding natural metric is $g=\left({\cdot},\,{\cdot}\right)_{\mathcal{A}}$.
A.2 Nonlinear problems
For general nonlinear problems, there is no canonical choice of metric as in the linear case. In some instances, the structure of the problem suggests multiple metrics; the choice of the metric impacts the convergence and the computational cost of the optimization schemes. Examples include Variational Monte Carlo (VMC) [68, 42, 65] or Sobolev gradient flows [41, 19]. Below we outline some general strategies that can be used in nonlinear problems and are employed in our numerical experiments.
Newton’s method
Setting A.4.
Let $(\mathrm{V},a)$ be a Riemannian manifold and $f:\mathrm{V}\to\mathbb{R}$ a smooth function. We aim to solve $\inf_{u\in\mathrm{V}}E(u)$.
If $\mathrm{Hess}^{a}E(u)$ is symmetric positive definite for all $u\in\mathrm{V}$, a natural metric leading to Riemannian Newton’s method [15] is
${g_{u}}\left({v},\,{w}\right)=\mathrm{Hess}^{a}E(u)[v,w].$ (31)
In the case $(\mathrm{V},a)=(\mathrm{V},\left({\cdot},\,{\cdot}\right)_{\mathrm{V}})$ for a Hilbert space $\mathrm{V}$, this yields the Energy Natural Gradient Descent (ENGD) method [64, 65]. Note that to interpret ENGD as a Natural Gradient Descent, it is necessary that $\mathrm{Hess}^{a}E(u)$ is SPD for all $u\in\mathrm{V}$, as otherwise (31) does not define a valid metric.
Nonlinear least-squares problem: Gauss-Newton
Setting A.5.
Let $\mathrm{V},\mathrm{W}$ be Hilbert spaces, and let $\mathcal{R}:\mathrm{V}\to\mathrm{W}$ be a (possibly nonlinear) residual. We aim at finding $u\in\mathrm{V}$ such that $\mathcal{R}(u)=0$.
A natural approach is to minimize the nonlinear least-squares functional
$\inf_{u\in\mathrm{V}}E(u)\coloneqq\inf_{u\in\mathrm{V}}\frac{1}{2}\left\|{\mathcal{R}(u)}\right\|_{\mathrm{W}}^{2}.$ (32)
The Hessian of $f$ reads
$\mathrm{D}^{2}E(u)[v,w]=\left({\mathrm{D}\mathcal{R}(u)[v]},\,{\mathrm{D}\mathcal{R}(u)[w]}\right)_{\mathrm{W}}+\left({\mathcal{R}(u)},\,{\mathrm{D}^{2}\mathcal{R}(u)[v,w]}\right)_{\mathrm{W}},$
and a symmetric positive semidefinite approximation is given by the Gauss-Newton method [83]
${g_{u}}\left({v},\,{w}\right)=\left({\mathrm{D}\mathcal{R}(u)[v]},\,{\mathrm{D}\mathcal{R}(u)[w]}\right)_{\mathrm{W}}.$ (33)
This choice of the metric leads the Gauss-Newton NGD method [46, 39]. Note that $g_{u}$ is, in general, only semidefinite. While this does not pose numerical issues when using the pseudoinverse or damping, the positive definitness of $g_{u}$ is necessary to interpret it as a Riemannian metric. Whether this condition holds depends on the structure of the underlying nonlinear PDE.
Appendix B Modification of jaxopt.BFGS
We refer to the JAXopt implementation of the BFGS algorithm in https://github.com/google/jaxopt/blob/main/jaxopt/_src/bfgs.py (commit c818497). Using the notation from the implementation, lines 240-250 compute the BFGS inverse Hessian update [83] as:
$\mathbf{H}_{k+1}=(\mathbf{I}-\rho\mathbf{s}\mathbf{y}^{\top})\mathbf{H}_{k}(\mathbf{I}-\rho\mathbf{y}\mathbf{s}^{\top})+\rho\mathbf{s}\mathbf{s}^{\top},\quad\rho=\frac{1}{\mathbf{s}^{\top}\mathbf{y}}.$
This involves outer products $\mathbf{s}\mathbf{y}^{\top}$ and $\mathbf{s}\mathbf{s}^{\top}$ and two matrix-matrix multiplications. We replace it with a more efficient formulation that avoids full matrix-matrix products, using vector-matrix operations and rank-one updates:
$\mathbf{H}_{k+1}=\mathbf{H}_{k}+\left[\rho+\rho^{2}\mathbf{y}^{\top}(\mathbf{H}_{k}\mathbf{y})\right]\mathbf{s}\mathbf{s}^{\top}-\rho\left[(\mathbf{H}_{k}\mathbf{y})\mathbf{s}^{\top}+\mathbf{s}(\mathbf{H}_{k}\mathbf{y})^{\top}\right].$
This reformulation eliminates matrix-matrix multiplications, resulting in significant speedups while maintaining numerical stability and equivalence to the original update.
Appendix C Sensitivity analysis and hyperparameter tuning
We conduct a sensitivity analysis of NyströmNGD with respect to its key hyperparameters, using the 3D Poisson problem discretized via PINNs as a model problem. Specifically, we examine the impact of the regularization parameter $\mu$, the adaptivity of the Nyström rank parameter $\ell$, and the maximum number of iterations and tolerance for the CG algorithm. Notation follows Section˜4.2.
Regularization parameter $\mu$
We set $\mu=\gamma\cdot\epsilon_{\textrm{mach}}\cdot\hat{\lambda}_{1}$ to regularize the least squares problem, solving with $(\mathbf{G}(\boldsymbol{\theta})+\mu I)^{-1}$ instead of $\mathbf{G}(\boldsymbol{\theta})^{\dagger}$. We recommend $\gamma=p$, following the threshold for numerical rank used in standard software like NumPy and MATLAB. Setting $\gamma$ much larger than $p$ leads to excessive regularization and a search direction that deviates from the “exact” natural gradient, while setting it much smaller can introduce numerical instability. This behavior is confirmed by the results in the Table˜9. Also, lowering $\gamma$ increases the Nyström rank parameter $\ell$, raising computational cost, as reflected in the relative iteration times.
Table 9: Sensitivity analysis of the parameter $\gamma$. Final accuracy (relative $\mathrm{H}^{1}$ error), and iteration time relative to that with the default value $\gamma=p$.
$\boldsymbol{\gamma}$ Final accuracy Time (relative)
$\boldsymbol{p\times 100}$ $1.52\text{\times}{10}^{-04}$ $0.8$
$\boldsymbol{p\times 10}$ $9.21\text{\times}{10}^{-05}$ $0.9$
$\boldsymbol{p}$ $4.01\text{\times}{10}^{-05}$ $1.0$
$\boldsymbol{p\times 0.1}$ $70\text{\times}{10}^{-05} 1.1$1.1
$\boldsymbol{p\times 0.01}$ $3.91\text{\times}{10}^{-03}$ $1.3$
Rank adaptivity
We adapt the Nyström rank parameter $\ell$ as the smallest index satisfying $\hat{\lambda}_{\ell}/\mu<r$, following the recommendation $r=10$ from [31]. Decreasing $r$ increases the rank $\ell$, improving the quality of the preconditioner at the cost of higher computational expense. This trade-off is clearly reflected in the results in Table˜10. We find that the default value $r=10$ achieves a good balance between accuracy and per-iteration efficiency.
Table 10: Sensitivity analysis of the parameter $r$. Final accuracy (relative $\mathrm{H}^{1}$ error), and iteration time relative to that with the default value $r=10$.
$\boldsymbol{r}$ Final accuracy Time (relative)
1000 $2.67\text{\times}{10}^{-04}$ $0.8$
100 $1.18\text{\times}{10}^{-04}$ $0.9$
10 $4.01\text{\times}{10}^{-05}$ $1.0$
1 $2.21\text{\times}{10}^{-05}$ $1.2$
0.1 $1.89\text{\times}{10}^{-05}$ $1.3$
Conjugate Gradient
We set the CG relative tolerance as $\mathtt{rel\_tol}=\min(\kappa,\lVert\nabla L(\theta)\rVert_{2})$ following the Trust-Region method, which preserves the local quadratic convergence of Newton’s method while also ensuring global convergence and reducing computation far from critical points [20]. To obtain steps closer to the “exact” NGD direction, one can decrease $\kappa$ and increase the maximum number of CG iterations $\mathtt{maxit}$. As shown in Table˜11, increasing $\mathtt{maxit}$ and lowering $\kappa$ slightly raises runtimes, though the main per-iteration cost remains dominated by computing the preconditioner. Final accuracy varies minimally, and the default values $\kappa=0.1$ and $\mathtt{maxit}=20$ provide a good balance between accuracy and efficiency.
Table 11: Sensitivity analysis of the CG parameters $\kappa$ and $\mathtt{maxit}$. Final accuracy (relative $\mathrm{H}^{1}$ error), and in parentheses the iteration time relative to that with the default values $\kappa=0.1$ and $\mathtt{maxit}=20$.
$\boldsymbol{\kappa}$ $\boldsymbol{\mathtt{maxit}}$ 2 10 20 50 100
$\boldsymbol{10^{-1}}$ $3.72\text{\times}{10}^{-05}$ ($1.0$) $4.01\text{\times}{10}^{-05}$ ($1.0$) $4.01\text{\times}{10}^{-05}$ ($1.0$) $4.01\text{\times}{10}^{-05}$ ($1.0$) $4.01\text{\times}{10}^{-05}$ ($1.0$)
$\boldsymbol{10^{-2}}$ $3.52\text{\times}{10}^{-05}$ ($1.1$) $4.12\text{\times}{10}^{-05}$ ($1.1$) $4.27\text{\times}{10}^{-05}$ ($1.1$) $4.27\text{\times}{10}^{-05}$ ($1.1$) $4.27\text{\times}{10}^{-05}$ ($1.1$)
$\boldsymbol{10^{-3}}$ $4.59\text{\times}{10}^{-05}$ ($1.1$) $5.65\text{\times}{10}^{-05}$ ($1.1$) $6.55\text{\times}{10}^{-05}$ ($1.1$) $6.55\text{\times}{10}^{-05}$ ($1.1$) $6.55\text{\times}{10}^{-05}$ ($1.1$)
$\boldsymbol{10^{-6}}$ $3.33\text{\times}{10}^{-05}$ ($1.0$) $1.48\text{\times}{10}^{-05}$ ($1.1$) $1.70\text{\times}{10}^{-05}$ ($1.1$) $1.85\text{\times}{10}^{-05}$ ($1.1$) $1.78\text{\times}{10}^{-05}$ ($1.1$)
$\boldsymbol{10^{-10}}$ $3.33\text{\times}{10}^{-05}$ ($1.0$) $1.43\text{\times}{10}^{-05}$ ($1.1$) $1.52\text{\times}{10}^{-05}$ ($1.1$) $2.00\text{\times}{10}^{-05}$ ($1.3$) $1.98\text{\times}{10}^{-05}$ ($1.3$)
Appendix D Alternative preconditioners
The matrix $(\mathbf{G}(\boldsymbol{\theta})+\mu I)$ that we aim to precondition in NGD is symmetric positive definite, dense, well-approximated by a low-rank matrix, and accessible efficiently through matrix-vector products, but not in $O(1)$ elementwise time. In this setting, to the best of our knowledge, randomized Nyström preconditioning represents the state of the art for constructing an efficient preconditioner. To address concerns about the suitability of the Nyström approximation, we also tested alternative low-rank preconditioners, namely Pivoted Partial Cholesky, with greedy pivot selection [43], uniformly random pivot selection, and RPCholesky pivot selection [17]. Results are reported in Table˜12.
As already noted in Remark˜4.5, the main limitation of greedy pivot selection and RPCholesky in our context is that they require access to the entire diagonal of $\mathbf{G}(\boldsymbol{\theta})$, which involves computing matvecs with all canonical basis vectors. This is reflected in the longer runtimes in Table˜12 and would be exacerbated for larger networks, whereas NyströmNGD scales efficiently, as shown in Section˜5.1.2. In contrast, selecting pivots uniformly at random produces a significantly less effective preconditioner. Moreover, as noted in [17, Property 2.1], Pivoted Partial Cholesky is mathematically equivalent to Nyström with sampling matrices corresponding to the pivots. This supports the use of Gaussian random vectors in Nyström in our setting: they achieve similar accuracy without requiring explicit pivot selection.
Table 12: Comparison of different preconditioners for matrix-free NGD. Final accuracy (relative $\mathrm{H}^{1}$ error), and iteration time relative to that of Nyström NGD.
Preconditioner Final accuracy Time (relative)
NyströmNGD $2.92\text{\times}{10}^{-05}$ 11
Greedy Pivoted Chol. $2.18\text{\times}{10}^{-05}$ $4.5$
RPCholesky $2.61\text{\times}{10}^{-05}$ $4.7$
Randomly Pivoted Chol. $1.22\text{\times}{10}^{-02}$ $0.62$